### v1

# Janus-faces of Temporal Constraint Languages: A dichotomy of expressivity

Abstract.
The Bodirsky-Kára classification of temporal constraint languages stands as one of the earliest and most seminal complexity classifications within infinite-domain Constraint Satisfaction Problems (CSPs), yet it remains one of the most mysterious in terms of algorithms and algebraic invariants for the tractable cases. We show that those temporal languages which do not pp-construct EVERYTHING (and thus by the classification are solvable in polynomial time) have, in fact, very limited expressive power as measured by the graphs and hypergraphs they can pp-interpret. This limitation yields many previously unknown algebraic consequences, while also providing new, uniform proofs for known invariance properties. In particular, we show that such temporal constraint languages admit $4$-ary pseudo-Siggers polymorphisms – a result that sustains the possibility that the existence of such polymorphisms extends to the much broader context of the Bodirsky-Pinsker conjecture.
When Ianus can’t express it all
He tries in vain, he hits a wall
When for $\mathbb{K}_{3}$ no way he knows
His face of pseudo-loops he shows
footnotetext: The Roman god Janus (Latin: Ianus), traditionally depicted with two faces looking in opposite directions, embodies the duality of temporal perspective. Funded by the European Union (ERC, POCOCOP, 101071674). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This research was funded in whole or in part by the Austrian Science Fund (FWF) [I5948]. For the purpose of Open Access, the authors have applied a $C$ BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission. The first author is a recipient of a DOC Fellowship of the Austrian Academy of Sciences at the Technische Universität Wien.
Johanna Brunar, Michael Pinsker, and Moritz Schöbi
1. Introduction
1.1. Constraint Satisfaction Problems
The Constraint Satisfaction Problem induced by a relational structure $\mathbb{A}$, denoted by $\operatorname{\textnormal{CSP}}(\mathbb{A})$, is the computational problem of deciding, given a finite input structure $\mathbb{B}$ of the same signature as $\mathbb{A}$, whether there exists a homomorphism from $\mathbb{B}$ to $\mathbb{A}$, i.e. a map preserving all relations. The underlying structure $\mathbb{A}$ will often be referred to as template structure. This notion of fixed-template CSPs provides a uniform framework for modelling many classical computational problems such as graph-colouring problems, $3$-SAT, or solving equations. In 2017, two independent confirmations of a complexity dichotomy for CSPs induced by arbitrary finite structures – conjectured already in [FV93, FV98] – marked a breakthrough in the research programme on the complexity of finite-domain CSPs: for every finite structure $\mathbb{A}$, either $\operatorname{\textnormal{CSP}}(\mathbb{A})$ is solvable in polynomial time, or it is NP-complete [Bul17, Zhu17, Zhu20a], contrasting Ladner’s theorem [Lad75]. Yet, prominent computational problems that can be phrased as the CSPs of a fixed template $\mathbb{A}$ require the domain of $\mathbb{A}$ to be of countably infinite size – for example, the problem of deciding whether a given finite digraph contains a cycle can be phrased as CSP over $(\mathbb{Q};<)$, but cannot be modelled in this way by any finite template. There provably does not exist a complexity dichotomy for CSPs with infinite templates, not even if the template is ‘close to finite’ in the sense of $\omega$-categoricity [BG08, GJK+20]. For certain countably infinite $\omega$-categorical structures, namely first-order reducts of finitely bounded homogeneous structures, though, a P/NP-complete complexity dichotomy has been conjectured by Bodirsky and Pinsker more than a decade ago (see [BPP21, BP16, BKO+17, BKO+19] for various formulations of the conjecture). The conjecture remains wide open in its generality, but has been verified for several classes of structures fitting into the conjectured framework, including temporal constraint languages [BK08b, BK10], equality constraint languages [BK08a], phylogeny CSPs [BJP16, BJP17], the universal homogeneous poset [KP17, KP18], MMSNP [BMM18, BMM21], first-order reducts of any homogeneous undirected graph [BMPP19] including the random graph [BP15a], the universal homogeneous tournament [MP22, MP24], and graph orientation problems with forbidden tournaments [BGP25, BM24, FP25].
One of the earliest complexity classifications within the scope of the Bodirsky-Pinsker conjecture (which predates the conjecture and most likely inspired it) was obtained for the class of CSPs induced by temporal constraint languages, i.e. structures of the form $(\mathbb{Q};R_{1},R_{2},\dots)$ where each $R_{i}$ has a first-order definition in $(\mathbb{Q};<)$ – the rational numbers with the dense linear order [BK08b, BK10]. To this day, temporal constraint languages remain an active subject of study. They give rise to a natural and important class of CSPs, appearing in prominent problems from artificial intelligence such as temporal and spatial reasoning, see e.g. [JD97, BJ03, KJJ03, DJ05, BJM+24]. Yet despite their practical importance, temporal CSPs remain among the most challenging and least understood templates within the scope of the Bodirsky–Pinsker conjecture. From an algorithmic perspective, they appear fundamentally different from most other classes for which the Bodirsky-Pinsker conjecture has been verified: the general procedure from [BM16], which relies on canonical polymorphisms [BPT13] to show polynomial-time solvability of an infinite-domain CSP by reducing the problem to the CSP of a finite template, is not applicable in the temporal case. Moreover, the descriptive complexity of temporal CSPs is known to differ vastly from the finite [BR22], and their algebraic properties – crucial to the success of the so-called algebraic approach to CSPs – are still largely unknown. While ongoing research aims to gain deeper insight into the cases that are solvable in polynomial time through new algorithmic approaches [Mot25], our contribution is to shed light on the algebraic aspects.
1.2. The algebraic approach
The complexity of finite-domain CSPs is determined by the ‘expressive power’ of the underlying structure $\mathbb{A}$, which is encoded in the set $\operatorname{\textnormal{Pol}}(\mathbb{A})$ of all compatible finitary operations – the polymorphism clone of $\mathbb{A}$. The study of the computational complexity of CSPs via the polymorphisms they have is what is now understood as the algebraic approach to CSPs, first developed in [Jea98, BJK05] and unified by [BOP18]. The finite-domain CSP dichotomy now takes the following form: as the only source of NP-completeness for $\operatorname{\textnormal{CSP}}(\mathbb{A})$ stands the ability of $\mathbb{A}$ to pp-construct EVERYTHING, i.e. every finite structure (see Section 2.1 for the definition of a pp-construction). For the purpose of this paper, we call a structure that pp-constructs EVERYTHING omni-expressive. Every finite structure that lacks this property gives rise to a CSP solvable in polynomial time, and its polymorphism clone satisfies certain symmetries. Here, we say that the polymorphism clone $\operatorname{\textnormal{Pol}}(\mathbb{A})$ of a structure $\mathbb{A}$ satisfies an identity, if the identity is witnessed by operations contained in $\operatorname{\textnormal{Pol}}(\mathbb{A})$ for all evaluations of their arguments. For example, $\operatorname{\textnormal{Pol}}(\mathbb{A})$ always contains
  • a $6$-ary polymorphism $s$ witnessing the $6$-ary Siggers identity [Sig10]
    $s(x,y,x,z,y,z)\approx s(y,x,z,x,z,y),$ (1)
  • a $4$-ary polymorphism $s$ witnessing the $4$-ary Siggers identity [KMM15]
    $s(a,r,e,a)\approx s(r,a,r,e),$ (2)
  • a $6$-ary polymorphism $o$ witnessing the Olšák identities [Olš17]
    $o(x,x,y,y,y,x)\approx o(x,y,x,y,x,y)\approx o(y,x,x,x,y,y),$ (3)
  • for some $k\geq 3$, a $k$-ary polymorphism $w$ witnessing the weak near unanimity (WNU) identities [MM08]
    $w(y,x,x,\dots,x)\approx w(x,y,x,\dots,x)\approx\dots\approx w(x,x,\dots,x,y),$ (4)
  • for some $k\geq 3$, a $k$-ary polymorphism $c$ witnessing the cyclic identity [BK12]
    $c(x_{1},x_{2},\dots,x_{k})\approx c(x_{2},\dots,x_{k},x_{1}).$ (5)
An algebraic approach to CSPs via polymorphisms is for countably infinite structures only possible in the setting of $\omega$-categoricity. Indeed, if $\mathbb{A}$ is $\omega$-categorical, then the complexity of $\operatorname{\textnormal{CSP}}(\mathbb{A})$ is again captured within the ‘local’ algebraic structure of $\operatorname{\textnormal{Pol}}(\mathbb{A})$ [BP15b]. Moreover, for first-order reducts of finitely bounded homogeneous structures $\mathbb{A}$, it is conjectured that omni-expressivity remains the only source of NP-completeness of $\operatorname{\textnormal{CSP}}(\mathbb{A})$ [BPP21, BOP18, BKO+17, BKO+19]. Naturally, the question arises whether the identities (1)-(5) have counterparts for $\omega$-categorical structures that are not omni-expressive. This was answered affirmatively in [BP20] for a pseudo-version of identity (1): every $\omega$-categorical structure that does not pp-construct EVERYTHING has polymorphisms $s,u,$ and $v$ witnessing the identity
$u\circ s(x,y,x,z,y,z)\approx v\circ s(y,x,z,x,z,y).$ (6)
The pseudo-versions of identities (2)–(5) are defined analogously, by composing each operation symbol in the original identities with a new unary operation symbol. The corresponding statements for the pseudo-versions of both the cyclic identities (5) and the weaker WNU identities (4) are known to be false in general (for temporal structures that are not omni-expressive and fail to admit pseudo-cyclic polymorphisms, see e.g. [Bod21, Proposition 12.9.1]; an example of an $\omega$-categorical not omni-expressive structure with no pseudo-WNU polymorphisms is given in [BBK+23, Theorem 4]). It is, however, worth noting that the condition of satisfying a WNU or a cyclic identity differs fundamentally from the condition of satisfying the identities (1), (2), or (3) in that the former is in fact defined by an infinite disjunction of $k$-ary formulae for every $k\geq 3$, while the latter is a single formula. To the best of the authors’ knowledge, no $\omega$-categorical structures are known that are not omni-expressive and do not admit $4$-ary pseudo-Siggers or pseudo-Olšák polymorphisms, leaving open the possibility that the existence of these polymorphisms may in fact characterise non-omni-expressivity for $\omega$-categorical structures – a question highlighted in [Bod21, Question $22$]. Several indications point towards a positive answer: it is known [BPP21, Proposition 6.6] that the pseudo-versions of all sets of identities that characterise non-omni-expressivity in the finite carry over in the case of $\omega$-categorical not omni-expressive structures that adhere to the standard-reduction from [BM16]. Temporal constraint languages and phylogeny CSPs are the only completely classified classes within the scope of the Bodirsky-Pinsker conjecture that do not conform to this reduction. As such, they currently represent the only known candidates for counterexamples within this range. Moreover, every conservative $\omega$-categorical structure that is not omni-expressive admits $4$-ary pseudo-Siggers polymorphisms [BKNP25]. Finally, in a somewhat different direction, the absence of pseudo-Olšák polymorphisms is known to imply NP-completeness [Mot25] for $\omega$-categorical structures (though here the NP-hardness is not known to stem from omni-expressivity).
Despite the early classification of temporal constraint satisfaction problems and substantial progress in the general theory of CSPs since then, the algebraic invariants underlying temporal templates that are not omni-expressive have remained poorly understood. Prior to this work, the only identities known to be satisfied in this setting were those inherited from general results – the existence of $6$-ary pseudo-Siggers polymorphisms [BP20] and, assuming $P\neq NP$, pseudo-Olšák polymorphisms [Mot25] – as well as the existence of pseudo-WNU polymorphisms of all arities $k\geq 3$ [BR22, Proposition 7.27]. In this paper, we provide a uniform framework that not only captures the existence of the aforementioned $6$-ary pseudo-Siggers, pseudo-Olšák, and pseudo-WNU polymorphisms, but also establish a whole new family of identities satisfied by all temporal constraint languages that are not omni-expressive. Among these, the existence of 4-ary pseudo-Siggers polymorphisms stands out as a previously unknown representative.
1.3. Loop lemmata
In the early stages of the systematic research programme on CSPs, those induced by finite graphs were among the first to be studied. Observe that the classical $k$-colouring problem coincides with $\operatorname{\textnormal{CSP}}(\mathbb{K}_{k})$, where $\mathbb{K}_{k}$ denotes the clique on $k$ vertices. Vastly generalising the NP-completeness of the $k$-colouring problem for all $k\geq 3$, Hell and Nešetřil showed that every undirected non-bipartite graph is either omni-expressive (and hence its CSP is NP-complete), or has a loop (in which case it is entirely inexpressive, and in particular, its CSP trivial) [HN90]. This result was later extended to certain digraphs:
Theorem 1.1 ([BKN09]).
Let $\mathbb{G}$ be a finite smooth digraph of algebraic length $1$. Either $\mathbb{G}$ is omni-expressive or $\mathbb{G}$ contains a loop.
As first observed in [Sig10], loops in digraphs correspond to algebraic invariants. Given a finite digraph $\mathbb{G}$ and an enumeration $(i_{1},j_{1}),(i_{2},j_{2}),\dots,(i_{m},j_{m})$ of its edges, the identity
$s(x_{i_{1}},\dots,x_{i_{m}})\approx s(x_{j_{1}},\dots,x_{j_{m}})$
is called the $\mathbb{G}$-loop condition; in particular, the $6$-ary Siggers identity (1) is the $\mathbb{K}_{3}$-loop condition, and the $4$-ary Siggers identity (2) is the the loop condition induced by the digraph
e{e}a{a}r{r}
By standard techniques, Theorem 1.1 implies that the polymorphism clone of every not omni-expressive finite structure satisfies, in particular, the Siggers loop conditions.
Corollary 1.2 ([KMM15]).
Let $\mathbb{A}$ be a finite relational structure that is not omni-expressive. If $\mathbb{G}$ is any finite smooth digraph of algebraic length $1$, then $\operatorname{\textnormal{Pol}}(\mathbb{A})$ satisfies the $\mathbb{G}$-loop condition.
An oligomorphic subgroup $\Omega$ of the automorphism group of a countable digraph $\mathbb{G}$ has, in particular, finitely many orbits in its action on the domain of $\mathbb{G}$. One is thus led to consider the finite quotient of $\mathbb{G}$ modulo $\Omega$, whose vertices are the $\Omega$-orbits, and whose edges are induced from $\mathbb{G}$. Observe that a loop in this finite digraph comes from an edge in $\mathbb{G}$ between two vertices belonging to the same $\Omega$-orbit – a so-called pseudo-loop modulo $\Omega$. The non-existence of pseudo-loops in $\mathbb{G}$ has been identified as a source of computational hardness for $\operatorname{\textnormal{CSP}}(\mathbb{G})$ in several cases [BP20, BBK+23, BKNP25]. Following this direction, we prove a variant of Theorem 1.1 for the case of digraphs $\mathbb{G}$ that are pp-interpretable in a temporal constraint language $\mathbb{A}$ that is not omni-expressive. Roughly, this means that $\mathbb{G}$ is expressible in primitive positive logic over $\mathbb{A}$. Since the structure $(\mathbb{Q};<)$ is $\omega$-categorical, its automorphism group is oligomorphic. The latter naturally acts on $\mathbb{A}$ by automorphisms, and in fact also on $\mathbb{G}$ by the definition of a (pp-)interpretation. Hence, we may regard $\operatorname{\textnormal{Aut}}((\mathbb{Q};<))$ also as an oligomorphic subgroup of $\operatorname{\textnormal{Aut}}(\mathbb{G})$.
Theorem 1.3.
Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive. If $\mathbb{G}$ is any smooth digraph that is pp-interpretable in $\mathbb{A}$ and has pseudo-algebraic length $1$ modulo $\operatorname{\textnormal{Aut}}((\mathbb{Q};<))$, then $\mathbb{G}$ contains a pseudo-loop modulo $\operatorname{\textnormal{Aut}}((\mathbb{Q};<))$.
Theorem 1.3 reveals the dichotomous nature of temporal constraint languages in terms of their expressivity. Namely, any such language is either omni-expressive or inexpressive based on the digraphs it can pp-interpret; in the latter case, the only pp-interpretable digraphs are those containing pseudo-loops. This loss of expressivity comes with a gain of algebraic invariants. The $\mathbb{G}$-pseudo-loop condition induced by a finite digraph $\mathbb{G}$ arises from the $\mathbb{G}$-loop condition by composing either side of the identity with unary function symbols $u,v$ as in Equation 6. From Theorem 1.3, we derive an infinite family of identities that is satisfied in every temporal constraint language that is not omni-expressive. Among these, we conclude the existence of $4$-ary and $6$-ary pseudo-Siggers polymorphisms.
Corollary 1.4.
Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive. If $\mathbb{G}$ is any finite smooth digraph of algebraic length $1$, then $\operatorname{\textnormal{Pol}}(\mathbb{A})$ satisfies the $\mathbb{G}$-pseudo-loop condition.
Our result confirms that the full range of loop conditions whose satisfaction is known to characterise non-omni-expressivity in the finite setting lifts, in their pseudo-versions, to all not omni-expressive temporal structures. This crucially rules out temporal constraint languages as counterexamples, reinforcing the broader validity of these characterisations.
As it turns out, sets of identities of the form as (3) and (4) are better suited for deriving further structural properties that may, in particular, be used in algorithms [BK09, BK14, Bul06, Zhu17, Zhu20a, MP22, MP24, KTV24, PRSS25]. They are formalised through the notion of a $\mathbb{T}$-loop condition for hypergraphs $\mathbb{T}$ of arity $n\geq 3$, defined as a set of n1n-1 identities, explained in detail e.g. in [GJP19]. The following loop lemma for finite symmetric hypergraphs of suitable arity implies – in a similar way to how 1.2 is derived from Theorem 1.1 – the existence of WNU polymorphisms in all not omni-expressive finite structures. Here, a loop in an $n$-ary relation $R$ is a constant tuple $(a.\dots,a)\in R$.
Theorem 1.5 (follows from [BK12], see also [Zhu20b]).
Let $\mathbb{T}$ be a finite symmetric hypergraph of arity $n$ such that $p\nmid n$ for all prime numbers $p\leq|T|$. Either $\mathbb{T}$ is omni-expressive, or $\mathbb{T}$ contains a loop.
Extending the approach used in the proof of Theorem 1.3 to higher arities, we proof a version of Theorem 1.5 for the case symmetric hypergraphs pp-interpretable in a not omni-expressive temporal template. Note that in contrast to the finite version, no number-theoretic restrictions on the arity of $\mathbb{T}$ apply.
Theorem 1.6.
Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive. If $\mathbb{T}$ is any symmetric hypergraph of arity $n\geq 3$ that is pp-interpretable in $\mathbb{A}$, then $\mathbb{T}$ contains a pseudo-loop modulo $\operatorname{\textnormal{Aut}}((\mathbb{Q};<))$.
The pseudo-loop condition induced by a hypergraph is again obtained by composing all identities of the corresponding loop-condition with unary function symbols. Since the hypergraphs inducing the Olšák- and the WNU-identities, respectively, are symmetric, the existence of pseudo-Olšák polymorphisms and pseudo-WNU polymorphisms of all arities $n\geq 3$ follows from Theorem 1.6 in the standard way.
Corollary 1.7.
Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive. If $\mathbb{T}$ is any finite symmetric hypergraph of arity $n\geq 3$, then $\operatorname{\textnormal{Pol}}(\mathbb{A})$ satisfies the $\mathbb{T}$-pseudo-loop condition.
2. Preliminaries
2.1. Model-theoretic
For $n\in\mathbb{N}$, we denote the set $\{1,\dots,n\}$ by $[n]$. For an $n$-tuple $a=(a_{1},\dots,a_{n})$, we write $a_{i}$ or $\textnormal{pr}_{i}(a)$ for its $i$-th coordinate. By $\ker(a)$ we denote the subset of $n]\times[n]$ consisting of all pairs $(i,j)$ for which $a_{i}=a_{j}$. A relation $R$ on a set $A$ is a subset $R\subseteq A^{n}$ for some $n\in\mathbb{N}$, which is referred to as the arity of $R$. For a subset $I$ of $[n]$, we denote by $\textnormal{pr}_{I}(R)$ the $|I|$-ary relation obtained by projecting $R$ to all of its coordinates $i$ with $i\in I$. We also write $R|_{B_{1}\times\dots\times B_{n}}$ for the relation $R\cap(B_{1}\times\dots\times B_{n})$, where $B_{i}\subseteq A$. For the purpose of this paper, we assume all relations considered to be non-empty. A (relational) structure is a tuple $\mathbb{A}=(A;\mathcal{R})$ consisting of a set $A$ and a finite family $\mathcal{R}$ of relations on $A$ indexed by relational symbols. For structures $\mathbb{A}=(A;\mathcal{R})$ and $\mathbb{B}=(B;\mathcal{R}^{\prime})$ indexed the same relational symbols, a homomorphism from $\mathbb{A}$ to $\mathbb{B}$ is a mapping $f:A\to B$ such that for every relation $R\in\mathcal{R}$ and every $(r_{1},\dots,r_{n})\in R$, the tuple $(f(r_{1}),\dots,f(r_{n}))$ is contained in the corresponding relation of $\mathcal{R}^{\prime}$. The structures $\mathbb{A}$ and $\mathbb{B}$ are homomorphically equivalent if there exists homomorphisms both from $\mathbb{A}$ to $\mathbb{B}$, and from $\mathbb{B}$ to $\mathbb{A}$. An automorphism of $\mathbb{A}$ is a bijective homomorphism $f:A\to A$ whose inverse mapping $f^{-1}$ is also a homomorphism. By $\operatorname{\textnormal{Aut}}(\mathbb{A})$ we denote the automorphism group of $\mathbb{A}$. In the following, by componentwise application we will understand all mappings also as functions on $n$-tuples.
A first-order formula $\phi$ is called primitive positive (pp) over a set consisting of relational symbols $R_{1},\dots,R_{n}$ if it involves only the predicates $R_{i}$, existential quantification, and conjunction. A relation is pp-definable in a structure $\mathbb{A}=(A;\mathcal{R})$ if it is definable by a pp-formula over the relational symbols from $\mathcal{R}$. A structure $\mathbb{B}$ is pp-definable in $\mathbb{A}$ if all of its relations are. We say that $\mathbb{A}$ $p$-interprets $\mathbb{B}$ if there exist $k\geq 1$ and a partial surjective map $h:A^{k}\to B$ with the property that for every relation $R\subseteq B^{n}$ that is either a relation of $\mathbb{B}$, the equality relation on $B$, or $B$, its preimage $h^{-1}(R)$ – regarded as a relation of arity $nk$ on $A$ – is pp-definable in $\mathbb{A}$. If $\mathbb{A}$ $p$-interprets $\mathbb{B}$, then $\operatorname{\textnormal{CSP}}(\mathbb{B})$ is log-space reducible to $\operatorname{\textnormal{CSP}}(\mathbb{A})$ [BJK05]. We say that $\mathbb{A}$ $p$-constructs $\mathbb{B}$ if $\mathbb{B}$ is homomorphically equivalent to a structure that is pp-interpretable in $\mathbb{A}$. As homomorphically equivalent structures have the same CSPs, this also implies a log-space reduction from $\operatorname{\textnormal{CSP}}(\mathbb{B})$ to $\operatorname{\textnormal{CSP}}(\mathbb{A})$. We say that $\mathbb{A}$ pp-constructs EVERYTHING if $\mathbb{A}$ pp-constructs every finite structure; we then call $\mathbb{A}$ omni-expressive. Clearly, the CSP induced by an omni-expressive structure is NP-complete. By the finite-domain CSP dichotomy [Bul17, Zhu17, Zhu20a], the CSP of a finite structure that is not omni-expressive is always solvable in polynomial time.
A permutation group $\Omega$ acting on a set $A$ induces an equivalence relation on $A$: two elements $a,b\in A$ are equivalent if there is $\alpha\in\Omega$ such that $\alpha(a)=b$. The corresponding equivalence classes are called the orbits of $\Omega\curvearrowright A$, or simply $\Omega$-orbits. By componentwise evaluation, $\Omega$ acts on $A^{k}$ for every $k\in\mathbb{N}.$ The group is oligomorphic if this action has finitely many orbits for every $k\in\mathbb{N}$. A countable structure $\mathbb{A}$ is $\omega$-categorical if $\operatorname{\textnormal{Aut}}(\mathbb{A})$ is oligomorphic.
2.2. Graph-theoretic
A digraph is a relational structure of the form $\mathbb{G}=(G;E)$, where $E$ is binary. $\mathbb{G}$ is called smooth if $\textnormal{pr}_{1}(E)=\textnormal{pr}_{2}(E)$. If $E$ is symmetric, $\mathbb{G}$ is also called a graph. For $m\in\mathbb{N}$, we denote by $E^{m}$ the binary relation containing all tuples $(a_{0},a_{m})\in G\times G$ for which there exist $a_{1},\dots,a_{m-1}$ such that $(a_{i-1},a_{i})\in E$ for all $i\leq m$. We write $E^{-1}$ for the relation $\{(y,x):(x,y)\in E\}$. An $E$-walk from $a_{0}$ to $a_{n}$ is a finite sequence $p=(a_{0}E_{1}a_{1}E_{2}a_{2}\dots E_{n}a_{n})$, where $E_{i}\in\{E,E^{-1}\}$ and $(a_{i-1},a_{i})\in E_{i}$ for all $i$. We say that $p$ is closed if $a_{0}=a_{n}$. The algebraic length of $p$ is the number of occurrences of $E$ minus the number of occurrences of $E^{-1}$. A digraph $\mathbb{G}=(G;E)$ is said to have algebraic length $1$ if there exists a closed $E$-walk of algebraic length $1$. A digraph $\mathbb{G}$ is weakly connected if for all distinct $a,b\in G$ there exists an $E$-walk from $a$ to $b$. A subset $W\subseteq G$ is called a weakly connected component of $\mathbb{G}$ if it is maximal with respect to inclusion such that the digraph $\mathbb{G}|_{W}\coloneqq(W;E|_{W\times W})$ is weakly connected. A fence in $E$ from $a_{0}$ to $a_{2n}$ is an $E$-walk of the form $(a_{0}\dots a_{2n})$ where $(a_{2i-2},a_{2i-1})\in E$ and $(a_{2i-1},a_{2i})\in E^{-1}$ for all $i\leq n$. The vertices $a_{0},a_{2},\dots,a_{2n}$ are called the lower tips of $p$. For $m\in\mathbb{N}$, a fence in $E^{m}$ is also called an $m$-fence in $E$. We say that $\mathbb{G}$ is linked if there exists $m\in\mathbb{N}$ such that for any $a,b\in\textnormal{pr}_{1}(E)$ there exists an $m$-fence in $E$ from $a$ to $b$. A finite smooth digraph is linked if and only if it is weakly connected and has algebraic length $1$ [BK12, Claim 3.8].
A hypergraph is any structure $\mathbb{T}=(T;R)$ containing only one relation $R$, where $R$ is not necessarily binary. Let $n$ be the arity of $R$, and let $\Sigma$ be a subgroup of the symmetric group $\textnormal{Sym}(n)$ on $n$ elements. We say that $\mathbb{T}$ is $\Sigma$-invariant if for every $a\in R$ and $\pi\in\Sigma$ also $(a_{\pi(1)},\dots,a_{\pi(n)})\in R$. A hypergraph is symmetric if it is $\textnormal{Sym}(n)$-invariant, cyclic if it is invariant with respect to the group of cyclic shifts, and $2$-transitive if it is $\Sigma$-invariant for some $2$-transitive group $\Sigma$. A loop is a constant tuple $(a,\dots,a)\in R$.
2.3. Algebraic
A polymorphism of arity $n\geq 1$ of a structure $\mathbb{A}=(A;\mathcal{R})$ is a mapping $f:A^{n}\to A$ that preserves every $R\in\mathcal{R}$, i.e.  for every $R\in\mathcal{R}$ and any $r_{1},\dots,r_{n}\in R$, the tuple $f(r_{1},\dots,r_{n})$ is contained in $R$. For an $\omega$-categorical structure $\mathbb{A}$, a relation $R$ is pp-definable in $\mathbb{A}$ if and only if it is preserved by all polymorphisms of $\mathbb{A}$ [BN06]. The polymorphism clone of $\mathbb{A}$ is the set of all polymorphism of $\mathbb{A}$, and it is denoted by $\operatorname{\textnormal{Pol}}(\mathbb{A})$. Observe that it indeed forms a clone in the sense of universal algebra, as it is closed under composition and contains all projections.
An identity is a formal abstract expression of the form $f\approx g,$ where $f$ and $g$ are terms over a common functional language. A clone $\mathcal{C}$ of operations on $A$ is said to satisfy a given identity $f(x_{1},\dots,x_{n})\approx g(y_{1},\dots,y_{n})$ if the function symbols appearing in the terms $f$ and $g$ can be interpreted as elements of $\mathcal{C}$, such that the equality $f(x_{1},\dots,x_{n})=g(y_{1},\dots,y_{n})$ holds for any evaluation of the variables in $A$. The satisfaction of a set of identities is defined as the simultaneous satisfaction of all identities by means of the same interpretation of function symbols. The pseudo-version of an identity $f\approx g$ is the identity $u\circ f\approx v\circ g$, where $u$ and $v$ are fresh unary function symbols.
Given a finite hypergraph $\mathbb{T}$ of some arity $n\geq 2$ and an enumeration $(x_{1,1},\dots,x_{n,1}),\dots,$ $(x_{1,m},\dots,x_{n,m})$ of its edges, the $\mathbb{T}$-loop condition is the (n1)(n-1)-element set consisting of the identities
$s(x_{1,1},\dots,x_{1,m})\approx s(x_{2,1},\dots,x_{2,m})\approx\dots\approx s(x_{n,1},\dots,x_{n,m}).$
Similarly, the $\mathbb{T}$-pseudo-loop condition is the set containing the pseudo-identities
$u_{1}\circ s(x_{1,1},\dots,x_{1,m})\approx u_{2}\circ s(x_{2,1},\dots,x_{2,m})\approx\dots\approx u_{n}\circ s(x_{n,1},\dots,x_{n,m}).$
2.4. Temporal constraint languages
A temporal relation is a relation first-order definable in $(\mathbb{Q};<)$. A temporal constraint language is a relational structure all of whose relations are temporal. Clearly, the natural action of $\operatorname{\textnormal{Aut}}(\mathbb{Q})\coloneqq\operatorname{\textnormal{Aut}}((\mathbb{Q};<))$ on $\mathbb{Q}^{k}$ has finitely many orbits for every $k\in\mathbb{N}$. The automorphisms of a structure preserve all relations first-order definable in it. As a consequence, the automorphism group of a first-order reduct always contains the original automorphisms. It follows that the first-order reducts of $\omega$-categorical structures, and in particular all temporal constraint languages, are themselves $\omega$-categorical.
Let $F$ be a set of operations on $\mathbb{Q}$. We say that the clone generated by $F$ is the smallest clone of operations $\mathcal{C}$ that contains $F\cup\operatorname{\textnormal{Aut}}(\mathbb{Q})$ and is closed under interpolation; that is, an operation $g$ belongs to $\mathcal{C}$ if and only if for every finite subset $A\subseteq\mathbb{Q}$, there is $f\in\mathcal{C}$ agreeing with $g$ on $A$. In the case that $F$ contains only one operation $f$, we say that $f$ generates $g$. The clones generated by an operation $f$ and a set $F$ of operations are denoted by $\langle f\rangle$ and $\langle F\rangle$, respectively. For $S\subseteq\mathbb{Q}^{k}$ and an operation $f$ on $\mathbb{Q}$, we denote by $\langle S\rangle_{f}$ the smallest subset of $\mathbb{Q}^{k}$ that contains $S$ and is preserved by every operation of $\langle f\rangle.$
The dual of an $n$-ary operation $f$ on $\mathbb{Q}$ is the operation $\operatorname{dual}(f)$ defined by
$\operatorname{dual}(f)(x_{1},\dots,x_{n})\coloneqq-f(-x_{1},\dots,-x_{n}).$
For a set $F$ of operations, $\operatorname{dual}(F)$ denotes the set containing all duals of operations in $F$. It is not hard to see that an operation $f$ preserves a relation $R$ if and only if its dual preserves the relation $R=\{(-a_{1},\dots,-a_{n}):(a_{1},\dots,a_{n})\in R\}$ [BK10]. Moreover, as the dual of an automorphism of $(\mathbb{Q};<)$ is again an automorphism, for every operation $f$ we have $\langle\operatorname{dual}(f)\rangle=\operatorname{dual}(\langle f\rangle)$.
The Bodirsky-Kára classification (Theorem 2.3) identifies all not omni-expressive temporal constraint languages by the existence of specific polymorphisms, whose definitions are provided below. Since a temporal relation is preserved by an operation $f$ if and only if it is preserved by every operation in the clone generated by $f$, we do not need to distinguish between operations that generate the same clone. Figure 1 provides an illustration of the classification.
Definition 2.1.
The binary operation $\min:\mathbb{Q}^{2}\to\mathbb{Q}$ maps two values $x$ and $y$ to the smaller of the two values. Let $\alpha,\beta,\gamma$ be any endomorphisms of $(\mathbb{Q};<)$ satisfying for all $\epsilon>0,x\in\mathbb{Q}$
$\alpha(x)<\beta(x)<\gamma(x)<\alpha(x+\epsilon).$
The two operations $\operatorname{mi}$ and $\operatorname{mx}$ are defined as
$\operatorname{mi}(x,y)\coloneqq\begin{cases}\alpha(\min(x,y))&\textnormal{if}\ x=y\\ \beta(\min(x,y))&\textnormal{if}\ x<y\\ \gamma(\min(x,y))&\textnormal{if}\ x>y,\end{cases}\textnormal{ and }\operatorname{mx}(x,y)\coloneqq\begin{cases}\alpha(\min(x,y))&\text{if }x\neq y,\\ \beta(\min(x,y))&\text{if }x=y.\end{cases}$
As explained in [BK10], such endomorphisms can be constructed inductively. In this construction, one can easily enforce both that any of the endomorphisms has finitely many specified fixed points, or none at all. Note that our definition of $\operatorname{mi}$ differs slightly from the original one in [BK10]. The two operations generate the same clone, and in particular each other. Our definition of $\operatorname{mi}$ is in line with [BR22].
Definition 2.2.
For $q\in\mathbb{Q}$, let ${\ell\ell}_{q}:\mathbb{Q}^{2}\to\mathbb{Q}$ be a binary operation satisfying ${\ell\ell}_{q}(x,y)\leq{\ell\ell}_{q}(x^{\prime},y^{\prime})$ if and only if one of the following cases applies:
  • $x\leq q\ \&\ x<x^{\prime}$
  • $x\leq q\ \&\ x=x^{\prime}\ \&\ y<y^{\prime}$
  • $x,x^{\prime}>q\ \&\ y<y^{\prime}$
  • $x,x^{\prime}>q\ \&\ y=y^{\prime}\ \&\ x<x^{\prime}$
Such an operation exists, is injective by definition, and all operations satisfying these conditions generate the same clone. Clearly, ${\ell\ell}_{q}$ generates ${\ell\ell}_{p}$ for $p\neq q$. By $\ell\ell$ we denote the operation ${\ell\ell}_{0}$.
Theorem 2.3 (Bodirsky-Kára classification [BK10]).
Let $\mathbb{A}$ be a temporal constraint language. If $\mathbb{A}$ is not omni-expressive, then it is preserved by one of $\min$, $\operatorname{mi}$, $\operatorname{mx}$, $\ell\ell$, their duals, or a constant operation, and $\operatorname{\textnormal{CSP}}(\mathbb{A})$ is solvable in polynomial time.
Figure 1. Polymorphisms of temporal constraint languages
[…]
For the proofs of Theorems 1.3 and 1.6, we will additionally draw on two binary operations $\operatorname{\textnormal{lex}}$ and $pp$. As $\ell\ell$ and $\operatorname{dual}(\ell\ell)$ generate $\operatorname{\textnormal{lex}}$, and each of $\min,\operatorname{mi},$ and $\operatorname{mx}$ generates $pp$, it follows that one of $\operatorname{\textnormal{lex}}$, $p$,$p$, and $\operatorname{dual}(pp)$ is contained in the polymorphism clone of every not omni-expressive constraint language. It is worth noting, however, that the presence of $\operatorname{\textnormal{lex}}$ or $pp$ in a polymorphism clone does not, on its own, imply polynomial-time solvability of the underlying template. For example, the temporal constraint language modelling the classical Betweenness problem is preserved by $\operatorname{\textnormal{lex}}$, yet is NP-complete [Opa79].
Definition 2.4.
Let $\operatorname{\textnormal{lex}}:\mathbb{Q}^{2}\to\mathbb{Q}$ be a binary operation satisfying $\operatorname{\textnormal{lex}}(x,y)\leq\operatorname{\textnormal{lex}}(x^{\prime},y^{\prime})$ if and only if
  • $x<x^{\prime}$ or
  • $x=x^{\prime}\ \&\ y<y^{\prime}$.
Clearly, such an operation exists. It is easy to see that all operations with these properties generate the same clone [BK10, Observation 1]. As $\ell\ell$ and $\operatorname{dual}(\ell\ell)$ also satisfy them when restricted to $\{(x,y)\in\mathbb{Q}\times\mathbb{Q}\colon x,y<0\}$ and $\{(x,y)\in\mathbb{Q}\times\mathbb{Q}\colon x,y>0\}$, respectively, both of them generate $\operatorname{\textnormal{lex}}$. Note that lex is, by definition, injective.
Definition 2.5.
For $q\in\mathbb{Q}$, let $pp_{q}:\mathbb{Q}^{2}\to\mathbb{Q}$ be a binary operation satisfying $pp_{q}(x,y)\leq pp_{q}(x^{\prime},y^{\prime})$ if and only if
  • $x\leq q$ and $x\leq x^{\prime}$ or
  • $x,x^{\prime}>q$ and $y\leq y^{\prime}$.
Again, $pp_{q}$ generates $pp_{p}$ for $p\neq q$. By $pp$, we denote the operation $pp_{0}.$ It is generated from $\min,\operatorname{mx},$ and $\operatorname{mi}$ [BK10, Lemma 23].
For every $k\in\mathbb{N}$, we denote by $\sim_{k}$ the orbit-equivalence of $\operatorname{\textnormal{Aut}}(\mathbb{Q})\curvearrowright\mathbb{Q}^{k}$. In other words, for $x,y\in\mathbb{Q}^{k}$ we have $x\sim_{k}y$ if and only if $x_{i}\leq x_{j}\leftrightarrow y_{i}\leq y_{j}$ for all $i,j\leq k$. A pseudo-loop in an $n$-ary relation $R$ on $\mathbb{Q}^{k}$ is a tuple $(a_{1},\dots,a_{n})$ with $a_{1}\sim_{k}\dots\sim_{k}a_{n}.$ The $\sim_{k}$-factor of $R$ is the relation defined on the set of all $\sim_{k}$}-classes of $\mathbb{Q}^{k}$ that contains a tuple $(A_{1},\dots,A_{n})$ of $\sim_{k}$-classes if and only if there exists a tuple $(a_{1},\dots,a_{n})\in R$ such that $a_{i}\in A_{i}$ for every $i\in[n].$ We say that a binary relation $E$ on $\mathbb{Q}^{k}$ has pseudo-algebraic length $1$ if the $\sim_{k}$-factor of $E$ admits a closed walk of algebraic length $1$. Whenever convenient, we naturally consider $n$-tuples on $\mathbb{Q}^{k}$ as tuples on $\mathbb{Q}^{nk}.$
Finally, we note that if $\mathbb{G}$ is pp-interpretable in a temporal constraint language $\mathbb{A}$, then every polymorphism of $\mathbb{A}$ induces a polymorphism of $\mathbb{G}$ [BJK05]: this follows from the fact that we can regard $\mathbb{G}$ as a structure whose domain is a pp-definable subset of some power of $\mathbb{Q}$ factored by a pp-definable equivalence relation, with all its relations being pp-definable as well. In particular, this is true for the automorphisms of $\mathbb{Q}$, hence $\operatorname{\textnormal{Aut}}(\mathbb{Q})\curvearrowright\mathbb{G}$. If $\mathbb{G}=(G;E)$ is a digraph, a pseudo-loop in $\mathbb{G}$ modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ is an edge whose vertices are contained in the same orbit of $\operatorname{\textnormal{Aut}}(\mathbb{Q})\curvearrowright\mathbb{G}$. We say that $\mathbb{G}$ has pseudo-algebraic length $1$ modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ if there exists an $E$-walk of algebraic length $1$ so that its start point and its end point are contained in the same $\operatorname{\textnormal{Aut}}(\mathbb{Q})$-orbit. Similarly, a pseudo-loop modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ in an $n$-ary hypergraph $\mathbb{T}=(T;R)$ that is pp-interpretable in a temporal constraint language is an edge $(a_{1},\dots,a_{n})\in R$ such that all $a_{i}$ belong to the same $\operatorname{\textnormal{Aut}}(\mathbb{Q})$-orbit.
3. Chasing Orbits
For the proofs of Theorems 1.3 and 1.6, a pseudo-loop is constructed via a recursive procedure that, at each step, aligns the minimal elements of appropriately chosen tuples. We only consider relations on some power of $\mathbb{Q}$ that are preserved by one of $\min,\operatorname{mi},\operatorname{mx},$ or $\ell\ell$. The general case follows easily as discussed in the proof of Theorem 1.3.
Definition 3.1.
Let $k\geq 1$.
  • For every $a\in\mathbb{Q}^{k}$, we denote by $\min(a)$ the minimal entry of $a$. By $\operatorname{\textnormal{minx}}(a)$ we denote the set of all indices $i\leq k$ for which $a_{i}=\min(a)$.
  • For all $a,b\in\mathbb{Q}^{k}$ and $I\subseteq[k]$, we write $a\sim_{I}b$ if $\textnormal{pr}_{I}(a)\sim_{|I|}\textnormal{pr}_{I}(b)$.
  • For $1\leq m\leq k$ and $a\in\mathbb{Q}^{k}$, we set $I_{m}(a)$ for the set of those indices $i$ such that $a_{i}$ is among the $m$ smallest values that appear in $a$.
  • For $a,b\in\mathbb{Q}^{k}$ we write $a<b$ if $a_{i}<b_{i}$ for all $i\in[k].$
Given an $n$-ary relation $R$ on $\mathbb{Q}^{k}$, we iteratively find tuples $a^{m}\in R$ so that, for each $m\leq k$ and for all $i,j\in[n]$,
  1. (1)
    $I_{m}(a_{i}^{m})=I_{m}(a_{j}^{m})$, i.e. the $m$ smallest values of $a_{i}^{m}$ and $a_{j}^{m}$ are taken on the same set of coordinates;
  2. (2)
    $a^{m}_{i}\sim_{I_{m}(a^{m}_{i})}a^{m}_{j}$, i.e. the subtuples of $a^{m}_{i}$ and $a^{m}_{j}$ containing precisely the $m$ smallest values of $a^{m}_{i}$ and $a^{m}_{j}$, respectively, are contained in the same $\sim_{m}$-class;
  3. (3)
    $\textnormal{pr}_{I_{m}(a^{m+1}_{i})}(a^{m+1}_{i})\sim_{m}a^{m}_{i}$, i.e. the $\sim_{m}$-class of the subtuple of $a^{m+1}_{i}$ containing the $m$ smallest entries of $a^{m+1}_{i}$ coincides with the $\sim_{m}$-class of the subtuple of $a^{m}_{i}$ containing all of its $m$ smallest entries.
After at most $k$-many steps, this recursion yields a tuple all of whose coordinates lie in the same orbit of $\operatorname{\textnormal{Aut}}(\mathbb{Q})\curvearrowright\mathbb{Q}^{k}$, i.e. a pseudo-loop in $R$.
Definition 3.2.
Let $a$ be an $n$-tuple on $\mathbb{Q}^{k}$. By $M(a)$ we denote the set consisting of all $i\in[n]$ for which $\min(a_{i})=\min(a)$. We say that $a$ is min-clean if for all $i,j\in M(a)$ we have $\operatorname{\textnormal{minx}}(a_{i})=\operatorname{\textnormal{minx}}(a_{j})$.
The technical part of the proofs of Theorems 1.3 and 1.6 concerns the existence of min-clean tuples, and is deferred to Section 5. In the following, we construct a pseudo-loop under the assumption that min-clean tuples exist. We note an asymmetry between the cases distinguished by Theorem 2.3. When the template is preserved by $\ell\ell$, the existence of min-clean tuples only relies on the polymorphism $\operatorname{\textnormal{lex}}$ (which is generated by $\ell\ell$), while the subsequent construction uses the full strength of $\ell\ell$. In the other cases, the construction proceeds using $pp$ (which alone does not prevent omni-expressivity), and the existence of min-clean tuples is shown by employing the polymorphisms $\min,\operatorname{mi}$, and $\operatorname{mx}$, respectively. The construction is made up of a nested composition of $\ell\ell$ and $pp$, building on the following easy observation: let $a\in\mathbb{Q}^{k}$, $q\in\mathbb{Q}$, and define $J\coloneqq\{i\in[k]:a_{i}\leq q\}$. For every $b\in\mathbb{Q}^{k}$, the $\sim_{k}$-class of the $k$-tuple $pp_{q}(a,b)$ is determined only by the $\sim_{|J|}$-class of the $|J|$-tuple $\textnormal{pr}_{J}(a)$ and the $\sim_{(k-|J|)}$-class of the (k|J|)(k-|J|)-tuple $\textnormal{pr}_{[k]\setminus J}(b)$. The same holds true for the tuple ${\ell\ell}_{q}(a,b)$, provided that $\ker(a)=\ker(b)$.
For a binary operation $f$ and $m\geq 2$, we denote by $f^{[m]}$} the $m$-ary operation
${f^{[m]}}(x_{1},\dots,x_{m})\coloneqq f(x_{1},f(x_{2},\dots,f(x_{m-1},x_{m})\dots)).$
Note that $f^{[m]}$} is lies within every clone containing $f$.
3.1. Templates with pppp
Lemma 3.3.
Let $E$ be a binary smooth relation on $\mathbb{Q}^{k}$ that is preserved by $\langle pp\rangle$, and assume that $E$ contains a min-clean tuple $\binom{u^{1}}{v^{1}}$. Then there exist $I\subseteq k$, $m,n\geq 1,q\in\mathbb{Q}$, and $\binom{u^{2}}{v^{2}},\ldots,\binom{u^{n}}{v^{n}}\in E$ such that for all $u,v\in\mathbb{Q}^{k}$ the tuple defined by
$\binom{u^{\prime}}{v^{\prime}}\coloneqq{pp_{q}^{[n+1]}}\left(\binom{u^{1}}{v^{1}},\binom{u^{2}}{v^{2}},\dots,\binom{u^{n}}{v^{n}},\binom{u}{v}\right)$
satisfies:
  • $I_{m}(u^{\prime})=I_{m}(v^{\prime})=I$ and $u^{\prime}\sim_{I}v^{\prime}$;
  • $u^{\prime}\sim_{[k]\setminus I}u$ and $v^{\prime}\sim_{[k]\setminus I}v$.
Proof.
If $\min u^{1}=\min v^{1}$ then $\operatorname{\textnormal{minx}}u^{1}=\operatorname{\textnormal{minx}}v^{1}$. In this case, we set $I\coloneqq\operatorname{\textnormal{minx}}u^{1}$, $m\coloneqq 1$, $n\coloneqq 1$, and $q\coloneqq\min u^{1}$.
Otherwise, assume without loss of generality that $\min u^{1}<\min v^{1}$. We set $q\coloneqq\min u^{1}$, and inductively set $v^{i+1}\coloneqq u^{i}$, and pick $u^{i+1}$ arbitrarily so that $\binom{u^{i+1}}{v^{i+1}}\in T$. Set $J^{1}\coloneqq\operatorname{\textnormal{minx}}u^{1}$, and inductively $J^{i+1}\coloneqq\{j\;|\;u^{i+1}_{j}\leq q\}\setminus J^{1}\cup\cdots\cup J^{i}$. Finally, we pick $n\geq 1$ minimal such that $J^{n}$ is empty, set $I$ to be the union $J^{1}\cup\cdots\cup J^{n}$, and $m\coloneqq|I|$.
For all $i$ we have $u^{\prime}\sim_{J_{i}}u^{i}$ and $v^{\prime}\sim_{J_{i}}v^{i+1}=u^{i}$, hence $u^{\prime}\sim_{J_{i}}v^{\prime}$. Moreover $\textnormal{pr}_{J_{i}}(u^{\prime})<\textnormal{pr}_{J_{i+1}}(u^{\prime})$, and similarly $\textnormal{pr}_{J_{i}}(v^{\prime})<\textnormal{pr}_{J_{i+1}}(v^{\prime})$ for all $i$. It follows that $u^{\prime}\sim_{I}v^{\prime}$. Moreover, $\textnormal{pr}_{I}(u^{\prime})<\textnormal{pr}_{[k]\setminus I}(u^{\prime})$ and $\textnormal{pr}_{I}(v^{\prime})<\textnormal{pr}_{[k]\setminus I}(v^{\prime})$, hence $I_{m}(u^{\prime})=I_{m}(v^{\prime})=I$. Finally, it is easy to verify that $u^{\prime}\sim_{[k]\setminus I}u$ and $v^{\prime}\sim_{[k]\setminus I}v$. ∎
A similar statement holds true for higher-arity relations $R$ that are symmetric. In fact, provided the existence of min-clean tuples in $R$, it suffices that $R$ is cyclic.
Lemma 3.4.
Let $R$ be an $n$-ary cyclic relation on $\mathbb{Q}^{k}$ that is preserved by $\langle pp\rangle$, and assume that $R$ contains a min-clean tuple $u^{1}=(u^{1}_{1}\dots,u^{1}_{n})$. Then there exist $I\subseteq[k],q\in\mathbb{Q}$, and $u^{2},\dots,u^{n}\in R$ such that for all $u=(u_{1},\dots,u_{n})\in R$ the tuple $u^{\prime}=(u^{\prime}_{1},\dots,u^{\prime}_{n})$ defined by
$u^{\prime}\coloneqq{pp_{q}^{[n+1]}}\left(u^{1},u^{2},\dots,u^{n},u\right)$
satisfies:
  • $I_{1}(u^{\prime}_{i})=I$ for all $i\leq n$;
  • $u^{\prime}_{i}\sim_{[k]\setminus I}u_{i}$ for all $i\leq n$.
Proof.
Without loss of generality assume that $1\in M(u^{1})$, and set $q\coloneqq\min(u^{1}_{1})$. By min-cleanliness of $u^{1}$, its cyclic permutations $u^{2}\coloneqq(u^{1}_{n},u^{1}_{1},u^{1}_{2},\dots),\dots,u^{n}\coloneqq(u^{1}_{2},\dots,u^{1}_{1})$ and $I\coloneqq\operatorname{\textnormal{minx}}(u^{1}_{1})$ now have the required properties. ∎
We now get a pseudo-loop by inductively shrinking $k$. To ensure the existence of min-clean tuples in cyclic relations $R$ of arity $n\geq 3$, as shown in Section 5, we additionally require that $R$ be $\Sigma$-invariant for some $2$-transitive subgroup $\Sigma$ of $\textnormal{Sym}(n)$.
Lemma 3.5.
Let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by either $\langle\min\rangle$, $\langle\operatorname{mi}\rangle$, or $\langle\operatorname{mx}\rangle$. If $n=2$ and $R$ is smooth and of pseudo-algebraic length $1$, or $n\geq 3$ and $R$ is cyclic and $2$-transitive, then $R$ contains a pseudo-loop.
Proof.
We use induction on $k$. The claim is trivial for $k=1$. Assume we have already shown it for all $j<k$. Lemmata 5.1, 5.5, 5.9, 5.2, 5.7, and 5.10 provide min-clean tuples in all cases, and the fact that $pp$ is contained in all three clones in question allows us to use Lemmata 3.3 and 3.4, respectively. We prove the statement for the binary case; the case $n\geq 3$ follows by analogy. Let $\binom{u^{1}}{v^{1}}\in R$ be min-clean, and take $I,m,n,q,$ and $\binom{u^{2}}{v^{2}},\ldots,\binom{u^{n}}{v^{n}}\in R$ as in Lemma 3.3. By induction hypothesis applied to the relation $\textnormal{pr}_{[k]\setminus I}R$, there exists $\binom{u}{v}\in R$ such that $u\sim_{[k]\setminus I}v$. For $\binom{u^{\prime}}{v^{\prime}}$ defined as in Lemma 3.3, we then have:
  • $I_{m}(u^{\prime})=I_{m}(v^{\prime})=I$ and $u^{\prime}\sim_{I}v^{\prime}$;
  • $u^{\prime}\;\sim_{[k]\setminus I}\;u\sim_{[k]\setminus I}\;v\;\sim_{[k]\setminus I}\;v^{\prime}$.
Hence, $\binom{u^{\prime}}{v^{\prime}}$ is a pseudo-loop. ∎
3.2. Templates with $\ell\ell$
In order to adapt Lemma 3.3 to the setting of digraphs preserved by $\ell\ell$, we need to restrict to tuples that are ‘as injective as possible’.
Definition 3.6.
Let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$. For $S\subseteq R$, we set
$S^{\prime}\coloneqq\{t\in\langle S\rangle_{\operatorname{\textnormal{lex}}}\;|\;\ker t=\bigcap_{s\in S}\ker s\}.$
The relation $S^{\prime}$ inherits desirable properties from $S$:
Lemma 3.7.
If $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{\textnormal{lex}}\rangle$. If $S\subseteq R$ is non-empty, then $S^{\prime}\neq\varnothing$. If $n=2$ and $S$ is smooth, then so is $S^{\prime}$. If $\Sigma$ is a subgroup of $\textnormal{Sym}(n)$ and $S$ is $\Sigma$-invariant, then so is $S^{\prime}$.
Proof.
The first statement is obtained by applying $\operatorname{\textnormal{lex}}^{[m]}$} of a sufficiently high arity $m$ to tuples in $S$ until the kernel of the resulting tuple can no more be refined. Suppose that $n=2$ and $S$ is smooth. To see that $S^{\prime}$ is smooth, pick $t\in S^{\prime}$, $f\in\langle\operatorname{\textnormal{lex}}\rangle$ of some arity $m\in\mathbb{N}$, and $s^{1},\dots,s^{m}\in S$ such that $t=f(s^{1},\dots,s^{m}).$ We show that $t_{1}\in\textnormal{pr}_{2}(S^{\prime})$, the converse statement for $t_{2}$ is shown analogously. By smoothness of $S$, for every $i\leq m$ there exists $r^{i}\in\mathbb{Q}^{k}$ such that $\binom{r^{i}}{s^{i}_{1}}\in S$. Let $w\coloneqq f(r^{1},\dots,r^{m})$. Picking any $\binom{u}{v}\in S^{\prime}$, we now have that $\operatorname{\textnormal{lex}}(\binom{w}{t_{1}},\binom{u}{v})\in S^{\prime}$. From $t_{1}\sim_{k}\operatorname{\textnormal{lex}}(t_{1},v)$, the statement follows. The last statement is clear from the definition of $S^{\prime}$. ∎
The following is an adaptation of Lemma 3.3 to the setting of smooth binary relations preserved by $\langle\ell\ell\rangle$.
Lemma 3.8.
Let $E$ be a binary smooth relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\ell\ell\rangle$. Let $S\subseteq E$ be smooth such that $S^{\prime}$ contains a min-clean tuple $\binom{u^{1}}{v^{1}}$. Then there exist $I\subseteq k$, $m,n\geq 1,q\in\mathbb{Q}$, and $\binom{u^{2}}{v^{2}},\ldots,\binom{u^{n}}{v^{n}}\in S^{\prime}$ such that for all $\binom{u}{v}\in S^{\prime}$ the tuple defined by
$\binom{u^{\prime}}{v^{\prime}}\coloneqq{\ell\ell_{q}^{[n+1]}}\left(\binom{u^{1}}{v^{1}},\binom{u^{2}}{v^{2}},\dots,\binom{u^{n}}{v^{n}},\binom{u}{v}\right)$
satisfies:
  • $I_{m}(u^{\prime})=I_{m}(v^{\prime})=I$ and $u^{\prime}\sim_{I}v^{\prime}$;
  • $u^{\prime}\sim_{[k]\setminus I}u$ and $v^{\prime}\sim_{[k]\setminus I}v$.
Proof.
We repeat the construction in the proof of Lemma 3.3, exchanging $pp_{q}$ for $\ell\ell_{q}$ and $E$ for $S^{\prime}$. The latter is possible because Lemma 3.7 guarantees smoothness of $S^{\prime}$, allowing us to choose suitable elements $v^{i+1}$. By definition of $S^{\prime}$, the tuple $(u^{\prime},v^{\prime})$ then has the required properties. ∎
Again, the existence of min-clean tuples as provided by Lemma 5.13 directly implies the existence of a pseudo-loop.
Lemma 3.9.
Let $E$ be a binary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\ell\ell\rangle$. If $S\subseteq E$ is smooth, has pseudo-algebraic length $1$, and $\textnormal{pr}_{1}(S)$ is weakly connected, then $S^{\prime}$ contains a pseudo-loop.
Proof.
We repeat the induction in the proof of Lemma 3.5. This time, Lemma 5.13 provides a min-clean tuple within $S^{\prime}$, and we use Lemma 3.8 instead of Lemma 3.3 to finish the proof. ∎
We proceed with the corresponding statement for hypergraphs preserved by $\langle\ell\ell\rangle$ that are $\Sigma$-invariant for some $2$-transitive cyclic group $\Sigma$.
Lemma 3.10.
Let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\ell\ell\rangle$. If $n\geq 3$ and $R$ is cyclic and $2$-transitive, then $R^{\prime}$ contains a pseudo-loop.
Proof.
Once more, we proceed by induction on $k$, employing Lemma 5.15 for the existence of a min-clean tuple within $R^{\prime}$, and the slight adaption of Lemma 3.4 to the setting of $\ell\ell$ changing $R$ to $R^{\prime}$. ∎
3.3. Proof of the main theorems
With the classification of not omni-expressive temporal constraint languages in hand, the remaining task is to combine the results established thus far to prove our main theorems.
See 1.3
Proof.
Replacing $\mathbb{G}$ by the preimage $\bar{\mathbb{G}}$ of $\mathbb{G}$ under the map of the pp-interpretation in $\mathbb{A}$, we may assume that $\mathbb{G}$ is a digraph defined on a finite power of $\mathbb{Q}$ whose edge relation $E$ is preserved by every polymorphism of $\mathbb{A}$. Indeed, by definition of a pp-interpretation, the preimage $\bar{E}$ of $E$ is pp-definable in $\mathbb{A}$, and therefore preserved by $\operatorname{\textnormal{Pol}}(\mathbb{A})$ [BN06]. A pseudo-loop in $\bar{E}$ yields a pseudo-loop in $E$ modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$. Moreover, $\bar{E}$ is smooth and has pseudo-algebraic length modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$.
By Theorem 2.3, $\mathbb{A}$ is preserved by one of $\ell\ell$, $\min$, $\operatorname{mi}$, $\operatorname{mx}$, their duals, or a constant operation. Accordingly, $\mathbb{G}$ is preserved by one of these operations, and thus by all operations in its generated clone. Clearly, every non-empty relation preserved by a constant operation contains even a loop. Moreover, $\mathbb{G}$ contains a pseudo-loop if and only if $\mathbb{G}$ does. Recalling that, for $f\in\{\ell\ell,\min,\operatorname{mi},\operatorname{mx}\}$, $\mathbb{G}$ is preserved by $\langle\operatorname{dual}(f)\rangle$ if and only if $\mathbb{G}$ is preserved by $\langle f\rangle$, this allows us to restrict ourselves to the cases where $\mathbb{G}$ is preserved by one of $\langle\ell\ell\rangle$, $\langle\min\rangle$, $\langle\operatorname{mi}\rangle$, and $\langle\operatorname{mx}\rangle$. In the first case, let $W$ be a weakly connected component of $\mathbb{G}$ of pseudo-algebraic length $1$ modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$, and set $S\coloneqq E|_{W\times W}$. By Lemma 3.9, $S^{\prime}\subseteq E$ contains a pseudo-loop modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$. In all the other cases, we can directly apply Lemma 3.5, yielding a pseudo-loop modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ within $E$. ∎
Similarly, putting together Lemmata 3.10 and 3.5, we obtain the corresponding pseudo-loop lemma for symmetric hypergraphs $\mathbb{T}$. By what we have shown, it in fact suffices for $\mathbb{T}$ to be cyclic and $2$-transitive. This amounts to a strictly weaker condition than full symmetry: for example, for every odd number $n\geq 7$, the alternating group $A_{n}$ is a proper subgroup of $\textnormal{Sym}(n)$ with these properties. It might be of interest to point out that pseudo-loop lemmata involving symmetry conditions provided by non-trivial group actions have also been established for finite hypergraphs (see [BK12, Bru23]). See 1.6
4. Pseudo-loop conditions from pseudo-loops
Following the standard method – established first by Siggers [Sig10] for finite structures, and adapted to $\omega$-categorical ones by Barto-Pinsker [BP20] – the existence of pseudo-loops in Theorems 1.3 and 1.6 implies the validity of the corresponding pseudo-loop conditions in all temporal constraint languages that are not omni-expressive. See 1.4
Proof.
Suppose that $\mathbb{G}=(\{1,\dots,n\};E_{\mathbb{G}})$ is a finite smooth digraph of algebraic length $1$, and $u\circ s(x_{i_{1}},\dots,x_{i_{m}})\approx v\circ s(x_{j_{1}},\dots,x_{j_{m}})$ is the corresponding $\mathbb{G}$-pseudo-loop condition. We show that for every $k\geq 1$ and for every $a_{1},\dots,a_{n}\in\mathbb{Q}^{k}$ there exists an $m$-ary operation $s\in\operatorname{\textnormal{Pol}}(\mathbb{A})$ such that $s(a_{i_{1}},\dots,a_{i_{m}})$ and $s(a_{j_{1}},\dots,a_{j_{m}})$ are contained in the same $\operatorname{\textnormal{Aut}}(\mathbb{Q})$-orbit. A standard compactness argument (as provided, for example, in [BP20, Lemma 4.2]) using oligomorphicity of $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ then yields the ‘global’ validity in $\operatorname{\textnormal{Pol}}(\mathbb{A})$ of the desired pseudo-identity.
To this end, take $k\geq 1$, and let $a_{1},\dots,a_{n}$ be arbitrary $k$-tuples of elements of $\mathbb{Q}$. Consider the binary relation $R\subseteq\mathbb{Q}^{k}\times\mathbb{Q}^{k}$ that consists precisely of all the tuples $(a_{i},a_{j})$ for which $(i,j)\in E_{\mathbb{G}}$, and let $E_{\mathbb{H}}\coloneqq\langle R\rangle_{\operatorname{\textnormal{Pol}}(\mathbb{A})}$. Since $E_{\mathbb{H}}$ is, by construction, preserved by all polymorphisms of $\mathbb{A}$, it is pp-definable in $\mathbb{A}$ [BN06]. It follows that $\mathbb{H}\coloneqq(A^{k};E_{\mathbb{H}})$ is pp-interpretable in $\mathbb{A}$. Moreover, observe that $\mathbb{H}$ is a smooth digraph of algebraic length $1$. Applying Theorem 1.3, we get a pseudo-loop modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ in $\mathbb{H}$. Unravelling definitions, this means that there exists $s\in\operatorname{\textnormal{Pol}}(\mathbb{A})$ such that $s(a_{i_{1}},\dots,a_{i_{m}})$ and $s(a_{j_{1}},\dots,a_{j_{m}})$ are contained in the same $\operatorname{\textnormal{Aut}}(\mathbb{Q})$-orbit. ∎
The corresponding statement for loop conditions induced by symmetric hypergraphs is derived from Theorem 1.6 in a similar way. Again, we may replace the word ‘symmetric’ by ‘cyclic and $2$-transitive’. See 1.7
5. Min-clean tuples
Recall the definition of a min-clean tuple: for an $n$-ary relation $R$ on $\mathbb{Q}^{k}$, we say that $t\in R$ is min-clean if $\operatorname{\textnormal{minx}}(t_{i})=\operatorname{\textnormal{minx}}(t_{j})$ for all $i,j\in M(t)$, where
$M(t)\coloneqq\left\{i\in[n]\colon\min(t_{i})=\min(t)\right\}.$
Moreover, a tuple $t\in R$ is min-ready in $R$ if the set $M(t)$ is minimal with respect to inclusion amongst all other sets $M(s)$ for $s\in R$. Clearly, every tuple $t\in R$ that satisfies $|M(t)|=1$ is both min-clean and min-ready in $R$.
5.1. Min-clean tuples for $\min$
Lemma 5.1.
Let $E$ be a binary relation on $\mathbb{Q}^{k}$. If $E$ is smooth and preserved by $\min$, then $E$ contains a min-clean tuple.
Proof.
If $E$ contains a tuple $t$ with $|M(t)|=1,$ then $t$ is min-clean. Otherwise, take $W$ to be any weakly connected component of $E$. Observe that by connectivity of $E|_{W\times W}$, all vertices $a,b$ contained in $W$ share the same minimal entry. It follows that $\operatorname{\textnormal{minx}}(\min(a,b))=\operatorname{\textnormal{minx}}(a)\cup\operatorname{\textnormal{minx}}(b)$ for all $a,b\in W$. Pick $t^{1},\dots,t^{m}\in E|_{W\times W}$ whose $\sim_{2k}$-classes represent all $\sim_{2k}$-classes appearing in $E|_{W\times W}$. Clearly, this is possible as there are only finitely many $\sim_{2k}$-classes on $\mathbb{Q}^{2k}$. Observe that by smoothness of $E|_{W\times W}$, the set of $\sim_{k}$-classes appearing among $t_{1}^{1},\dots,t_{1}^{m}$ coincides with the set of $\sim_{k}$-classes of $t_{2}^{1},\dots,t_{2}^{m}$. Therefore, the tuple $t\coloneqq{\min^{[m]}}(t^{1},\dots,t^{m})$ is min-clean because it satisfies $\operatorname{\textnormal{minx}}(t_{1})=\bigcup_{a\in W}\operatorname{\textnormal{minx}}(a)=\operatorname{\textnormal{minx}}(t_{2})$. ∎
As it turns out, in the case of symmetric relations preserved by $\min$, our endeavour to prove the existence of min-clean tuples leads us to the immediate conclusion that every such relation must contain an actual loop. In fact, it even suffices that the relation is cyclic.
Lemma 5.2.
Let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$. If $R$ is cyclic and preserved by $\min$, then $R$ contains a loop.
Proof.
Take $t\in R$ arbitrarily, and let $t^{1},\dots,t^{n}$ denote all cyclic permutations of the tuple $t$. Since $R$ is cyclic, we have $t^{i}\in R$ for every $i\leq n$. By symmetry of $\min^{[n]}$}, the tuple defined by ${\min^{[n]}}(t^{1},\dots,t^{n})$ is a loop in $R$. ∎
5.2. Min-clean tuples for $\operatorname{mi}$
Lemma 5.3.
Let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\operatorname{mi}$, $t^{1}\in R$ min-ready, and $t^{2},\dots,t^{m}\in R$. If $i\in M(t^{1})\cap\dots\cap M(t^{m}),$ $\operatorname{\textnormal{minx}}(t^{1}_{i})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{i})\neq\varnothing$, $\min(t_{i}^{1})=\dots=\min(t_{i}^{m})$, and $j\in M(t^{1})$, then also $j\in M(t^{1})\cap\dots\cap M(t^{m})$ and $\operatorname{\textnormal{minx}}(t^{1}_{j})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{j})\neq\varnothing.$
Proof.
Let $c=\min(t_{i}^{1})$, and choose mappings $\alpha,\beta$ and $\gamma$ satisfying the requirements in 2.1 in a way such that $\alpha(c)=c$. Consider the tuple $u\coloneqq{\operatorname{mi}^{[m]}}(t^{1},\dots,t^{m})$. Note that $\operatorname{\textnormal{minx}}(u_{i})=\operatorname{\textnormal{minx}}(t_{i}^{1})\cap\dots\cap\operatorname{\textnormal{minx}}(t_{i}^{m})$, and that $\min(u_{i})=c$. By min-readiness of $t^{1}$, also $\min(u_{j})=c$. As $\beta(c),\gamma(c)>\alpha(c)=c$, we must therefore have $j\in M(t^{1})\cap\dots\cap M(t^{m})$ and $\operatorname{\textnormal{minx}}(t_{j}^{1})\cap\dots\cap\operatorname{\textnormal{minx}}(t_{j}^{m})\neq\varnothing$. ∎
Lemma 5.4.
Let $E$ be a binary relation on $\mathbb{Q}^{k}$. Suppose that $W$ is a weakly connected component of $E$ of pseudo-algebraic length $1$. If $E$ is smooth and preserved by $\operatorname{mi}$, then either $E$ contains a tuple $t$ with $|M(t)|=1$, or $\bigcap_{a\in W}\operatorname{\textnormal{minx}}(a)\neq\varnothing$.
Proof.
If $|M(t)|=2$ for all $t\in E$, then by connectivity of $E|_{W\times W}$ we have $\min(a)=\min(b)$ for all vertices $a,b$ appearing in $W$. Take $a,b\in W$ arbitrarily. Since $E$ has pseudo-algebraic length $1$, the $\sim_{k}$-factor of the relation $E|_{W\times W}$ is smooth and $m$-linked for some $m\in\mathbb{N}$. Therefore, there exist $b^{\prime}\sim_{k}b$ and an $m$-fence in $E$ from $a$ to $b^{\prime}$, i.e. an $E$-walk of the form
$(x_{1,0}Ex_{1,1}E\dots Ex_{1,m}=x_{2,m}E^{-1}x_{2,m-1}E^{-1}\dots E^{-1}x_{2,0}=x_{3,0}E\dots E^{-1}x_{n,1}E^{-1}x_{n,0}),$
where $a=x_{1,0}$ and $b^{\prime}=x_{n,0}$. Since $b^{\prime}\sim_{k}b$, we have $\operatorname{\textnormal{minx}}(b)=\operatorname{\textnormal{minx}}(b^{\prime})$. We claim that there is a coordinate on which every lower tip of the above $m$-fence admits its minimal entry, that is
$\operatorname{\textnormal{minx}}(a)\cap\bigcap_{\ell<n}\operatorname{\textnormal{minx}}(x_{\ell,0})\cap\operatorname{\textnormal{minx}}(b^{\prime})\neq\varnothing.$ (7)
Indeed, applying Lemma 5.3 to the tuples $\binom{x_{1,m-1}}{x_{1,m}}$ and $\binom{x_{2,m-1}}{x_{1,m}}$, we obtain $\operatorname{\textnormal{minx}}(x_{1,m-1})\cap\operatorname{\textnormal{minx}}(x_{2,m-1})\neq\varnothing$. Consequently, we also get $\operatorname{\textnormal{minx}}(x_{1,m-2})\cap\operatorname{\textnormal{minx}}(x_{2,m-2})\neq\varnothing$, and, after $m$ steps, $\operatorname{\textnormal{minx}}(a)\cap\operatorname{\textnormal{minx}}(x_{2,0})\neq\varnothing$. We now apply Lemma 5.3 to the edges $\binom{a}{x_{1,1}},\binom{x_{2,0}}{x_{2,1}}$ and $\binom{x_{2,0}}{x_{3,1}}$, and deduce that also $\operatorname{\textnormal{minx}}(x_{1,m})\cap\operatorname{\textnormal{minx}}(x_{3,m})\neq\varnothing$. Continuing like this, we ultimately see that Equation 7 holds true.
Since the $\sim_{k}$-factor of $W$ is finite, we can select finitely many elements $a_{1},a_{2},\dots,a_{\ell}\in W$ such that each $\sim_{k}$-class appearing in $W$ is represented by some $a_{i}$. Moreover, we may choose these elements so that for every $i<\ell$, there exists an $m$-fence in $E$ connecting $a_{i}$ to $a_{i+1}$. By concatenating these fences, we obtain an $m$-fence whose set of lower tips contains all the elements $a_{1},a_{2},\dots,a_{\ell}$. Repeating the argument in the previous paragraph, we see that $\bigcap_{i\leq\ell}\operatorname{\textnormal{minx}}(a_{i})\neq\varnothing$. By the choice of the $a_{i}$, this intersection is equal to $\bigcap_{a\in W}\operatorname{\textnormal{minx}}(a)$. ∎
Lemma 5.5.
Let $E$ be a binary relation on $\mathbb{Q}^{k}$. If $E$ is smooth, has pseudo-algebraic length 1, and is preserved by $\operatorname{mi}$, then $E$ contains a min-clean tuple.
Proof.
Any tuple $t\in E$ with $|M(t)|=1$ is min-clean. Suppose that such a tuple does not exist in $E$, and fix a weakly connected component $W$ of $E$ that has pseudo-algebraic length $1$. It follows that all tuples in $W$ share the same minimal entry $c$. Pick $t^{1},\dots,t^{m}$ such that every $\sim_{2k}$-class of a tuple from $E|_{W\times W}$ is represented by one of $t^{1},\dots,t^{m}$. By smoothness of $E|_{W\times W}$, every $\sim_{k}$-class appearing among $t^{1}_{1},\dots,t^{m}_{1}$ coincides with some $\sim_{k}$-class appearing among $t^{1}_{2},\dots,t^{m}_{2}$, and vice versa. In the definition of $\operatorname{mi}$, choose $\alpha,\beta$ and $\gamma$ such that $\alpha(c)=c$. By Lemma 5.4, we have $\bigcap_{i\in[m]}\operatorname{\textnormal{minx}}(t^{i}_{j})=\bigcap_{a\in W}\operatorname{\textnormal{minx}}(x)\neq\varnothing$ for $j=1,2.$ The tuple $t\coloneqq{\operatorname{mi}^{[m]}}(t^{1},\dots,t^{m})$ now satisfies $\min(t_{1})=\min(t_{2})=c$, and $\operatorname{\textnormal{minx}}(t_{1})=\bigcap_{a\in W}\operatorname{\textnormal{minx}}(a)=\operatorname{\textnormal{minx}}(t_{2})$. ∎
Lemma 5.6.
Let $R$ be an $n$-ary 22-transitive relation on $\mathbb{Q}^{k}$ that is preserved by $\operatorname{mi}$. Either there is $t\in R$ with $|M(t)|=1,$ or $|M(t)|=n$ for all $t\in R$.
Proof.
Suppose that $t\in R$ satisfies $|M(t)|<n,$ and without loss of generality assume that $1\in M(t)$. By $2$-transitivity of $R$, there are tuples $t^{2},\dots,t^{n}\in R$ such that $1\in M(t^{i})$, $i\notin M(t^{i})$, and $\min(t^{i})=\min(t)$ for all $2\leq i\leq n$. In 2.1, choose $\alpha$ in the definition of $\operatorname{mi}$ such that $\alpha(\min(t))=\min(t).$ It then follows that the tuple $s\coloneqq{\operatorname{mi}^{[n-1]}}(t^{2},\dots,t^{n})$ satisfies $\min(s)=\min(t)$, and $M(s)=\{1\}$. ∎
Lemma 5.7.
Let $R$ be an $n$-ary cyclic and $2$-transitive relation on $\mathbb{Q}^{k}$ with $n\geq 3$. If $R$ is preserved by $\operatorname{mi}$, then $R$ contains a min-clean tuple.
Proof.
Suppose $\min_{t\in R}|M(t)|>1$, and take $t\in R$ arbitrarily. By Lemma 5.6, $t$ is min-ready and satisfies $\min(t_{1})=\dots=\min(t_{n})=:c$. We claim that $\bigcap_{i\leq n}\operatorname{\textnormal{minx}}(t_{i})\neq\varnothing.$ Indeed, define $t^{1}\coloneqq t$, and take $t^{2},\dots,t^{n-1}\in R$ to be permutations of $t$ satisfying $t^{i}_{1}=t_{1}$ and $t^{i}_{2}=i+1$ for all $2\leq i\leq n-1$. This is possible by $2$-transitivity of $R$. Moreover, we may choose $t^{n}\in R$ that is a permutation of $t$ and satisfies $t^{n}_{2}=t_{1}$ and $t^{n}_{3}=t_{3}$. Since $\bigcap_{i\leq n-1}\operatorname{\textnormal{minx}}(t_{1}^{i})=\operatorname{\textnormal{minx}}(t_{1})$, Lemma 5.3 applied to the tuples $t^{1},\dots,t^{n-1}$ gives $\bigcap_{i\leq n-1}\operatorname{\textnormal{minx}}(t^{i}_{j})\neq\varnothing$ for all $j\geq 2$. As, by definition, it holds that $\operatorname{\textnormal{minx}}(t^{n}_{3})=\operatorname{\textnormal{minx}}(t^{1}_{3})$ and hence $\bigcap_{i\leq n}\operatorname{\textnormal{minx}}(t^{i}_{3})\neq\varnothing$, we may apply Lemma 5.3 again to the tuples $t^{1},\dots,t^{n}$. We obtain $\bigcap_{i\leq n}\operatorname{\textnormal{minx}}(t^{i}_{j})\neq\varnothing$ for all $j\leq n$. By the choice of $t^{i}$, this intersection is for $j=2$ equal to $\bigcap_{i\geq 2}\operatorname{\textnormal{minx}}(t_{i})$.
Let now $s^{1},\dots,s^{n}$ denote all cyclic permutations of the tuple $t$. Since $R$ is cyclic, we have $s^{i}\in R$ for every $i\leq n$. Choose $\alpha$ as in 2.1 such that $\alpha(c)=c,$ and set $s\coloneqq{\operatorname{mi}^{[n]}}(s^{1},\dots,s^{n}).$ Observe that $\min(s)=c$, and $\operatorname{\textnormal{minx}}(s_{i})=\bigcap_{j\leq n}\operatorname{\textnormal{minx}}(t_{j})$ for all $i\leq n$. In particular, $s$ is min-clean. ∎
5.3. Min-clean tuples for $\operatorname{mx}$
Lemma 5.8.
Let $E$ be a binary relation on $\mathbb{Q}^{k}$. If $E$ is preserved by $\operatorname{mx}$, then either $E$ contains a tuple $t$ with $|M(t)|=1$, or for all $t,t^{\prime}\in E$ with $\min(t_{i})=\min(t^{\prime}_{i})$ and $\operatorname{\textnormal{minx}}(t_{i})=\operatorname{\textnormal{minx}}(t^{\prime}_{i})$ for $i\in\{1,2\}$, also $\operatorname{\textnormal{minx}}(t_{j})=\operatorname{\textnormal{minx}}(t^{\prime}_{j})$ for $j\neq i$.
Proof.
Suppose that all tuples $t$ satisfy $|M(t)|=2$. Without loss of generality, let $t,t^{\prime}\in E$ be such that $c\coloneqq\min(t_{1})=\min(t^{\prime}_{1})$ and $\operatorname{\textnormal{minx}}(t_{1})=\operatorname{\textnormal{minx}}(t^{\prime}_{1})$. Observe that for all $a,b\in\mathbb{Q}^{k}$ with $\min(a)=\min(b)$ it holds that $\min(\operatorname{mx}(a,b))=\beta(\min(a))$ if and only if $\operatorname{\textnormal{minx}}(a)\triangle\operatorname{\textnormal{minx}}(b)=\varnothing$, where $\beta$ is the endomorphism in the definition of $\operatorname{mx}$. The tuple $s\coloneqq\operatorname{mx}(t,t^{\prime})$ now satisfies $\min(s_{1})=\beta(c)$. By assumption, we have $\min(s_{2})=\min(s_{1})=\beta(c)$, and thus indeed $\operatorname{\textnormal{minx}}(t_{2})\triangle\operatorname{\textnormal{minx}}(t_{2}^{\prime})=\varnothing$. ∎
Lemma 5.9.
Let $E$ be a binary relation on $\mathbb{Q}^{k}$. If $E$ has pseudo-algebraic length $1$ and is preserved by $\operatorname{mx}$, then $E$ contains a min-clean tuple.
Proof.
Any tuple with $|M(t)|=1$ is min-clean. Suppose that $|M(t)|=2$ for all $t\in E$. Take $W$ a weakly connected component of pseudo-algebraic length $1$. As the $\sim_{k}$-factor of the relation $E|_{W\times W}$ is finite, smooth, weakly connected, and of algebraic length $1$, it is $m$-linked for some $m\in\mathbb{N}$. We claim that every $t\in E|_{W\times W}$ is min-clean. Indeed, for all $t\in E|_{W\times W}$ there are $t^{\prime}_{2}\sim_{k}t_{2}$ and an $m$-fence in $E|_{W\times W}$ connecting $t_{1}$ and $t^{\prime}_{2}$. Similar to the proof of Lemma 5.4, repeatedly applying Lemma 5.8 instead of Lemma 5.3, it follows that $\operatorname{\textnormal{minx}}(t_{1})=\operatorname{\textnormal{minx}}(t_{2})$.
Lemma 5.10.
Let $R$ be an $n$-ary 22-transitive relation on $\mathbb{Q}^{k}$ with $n\geq 3$. If $R$ is preserved by $\operatorname{mx}$, then $R$ contains a min-clean tuple.
Proof.
Suppose that $t\in R$ is not min-clean. In particular, we must have $|M(t)|>1$. Without loss of generality, assume $1,2\in M(t)$ and $\operatorname{\textnormal{minx}}(t_{1})\neq\operatorname{\textnormal{minx}}(t_{2})$. Observe that $\operatorname{\textnormal{minx}}(\operatorname{mx}(t_{1},t_{2}))=\operatorname{\textnormal{minx}}(t_{1})\triangle\operatorname{\textnormal{minx}}(t_{2})$, and $\min(\operatorname{mx}(t_{1},t_{2}))=\alpha(\min(t))$. By $2$-transitivity of $R$, there is a tuple $t^{\prime}\in R$ for which $t^{\prime}_{1}=t_{2}$ and $t^{\prime}_{2}=t_{1}$. The tuple $s\coloneqq\operatorname{mx}(t,t^{\prime})$ is contained in $R$, and it satisfies $M(s)=\{1,2\}$, $\min(s)=\alpha(\min t)$, and $\operatorname{\textnormal{minx}}(\textnormal{pr}_{1}(s))=\operatorname{\textnormal{minx}}(\textnormal{pr}_{2}(s))$. ∎
5.4. Min-clean tuples for $\ell\ell$
Lemma 5.11.
Let $R$ be preserved by $\operatorname{\textnormal{lex}}$, $t^{1}\in R$ min-ready, and $t^{2},\dots,t^{m}\in R$. If $i\in M(t^{1})\cap\dots\cap M(t^{m}),$ $\operatorname{\textnormal{minx}}(t^{1}_{i})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{i})\neq\varnothing$, and $j\in M(t_{1})$, then also $j\in M(t^{2})\cap\dots\cap M(t^{m})$ and $\operatorname{\textnormal{minx}}(t^{1}_{j})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{j})\neq\varnothing.$
Proof.
Consider the tuple $t$ defined by $t\coloneqq{\operatorname{\textnormal{lex}}^{[m]}}(t^{1},\dots,t^{m})$. Observe that $\operatorname{\textnormal{minx}}(t_{i})=\operatorname{\textnormal{minx}}(t^{1}_{i})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{i})$ and $\min(t)={\operatorname{\textnormal{lex}}^{[m]}}(\min(t^{1}_{i}),\dots,\min(t^{m}_{i}))$. By min-readiness of $t^{1}$, we also have $\min(t_{i})=\min(t_{j}),$ which by injectivity of $\operatorname{\textnormal{lex}}^{[m]}$} implies $\operatorname{\textnormal{minx}}(t^{1}_{j})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{j})\neq\varnothing.$
Lemma 5.12.
Let $E$ be a binary relation on $\mathbb{Q}^{k}$. Suppose that $W$ is a weakly connected component of $E$ of pseudo-algebraic length $1$. If $E$ is smooth and preserved by $\operatorname{\textnormal{lex}}$, then either $E$ contains a tuple $t$ with $|M(t)|=1$, or $\bigcap_{a\in W}\operatorname{\textnormal{minx}}(a)\neq\varnothing$.
Proof.
Once more, we proceed as in the proof of Lemma 5.4. This time, we repeatedly use Lemma 5.11 instead of Lemma 5.3. ∎
Lemma 5.13.
Let $E$ be a binary relation on $\mathbb{Q}^{k}$ that is smooth and preserved by $\operatorname{\textnormal{lex}}$. If $S\subseteq E$ is smooth itself, has pseudo-algebraic length $1$, and $\textnormal{pr}_{1}(S)$ is weakly connected, then $S^{\prime}$ contains a min-clean tuple.
Proof.
By Lemma 3.7, $S^{\prime}$ is non-empty. If $E$ contains a tuple $t$ with $|M(t)|=1$, then for every $s\in S^{\prime}$ the tuple $\operatorname{\textnormal{lex}}(t,s)$ is min-clean and contained in $S^{\prime}$.
Suppose now that $|M(t)|=2$ for all $t\in E$. By connectedness, all tuples in $S$ share the same minimal entry $c$. Choose $t^{1},\dots,t^{m}\in S$ such that every $\sim_{2k}$-class of an element in $S$ coincides with the $\sim_{2k}$-class of one of the tuples among $t^{1},\dots,t^{m}$. Once more, by smoothness of $S$, the set of all $\sim_{k}$-classes of the tuples $t^{1}_{1},\dots,t^{m}_{1}$ coincides with the set of all $\sim_{k}$-classes of the tuples $t^{1}_{2},\dots,t^{m}_{2}$. Thus, by the choice of $t^{1},\dots,t^{m}$, we have
$\bigcap_{i\in[m]}\operatorname{\textnormal{minx}}(t_{1}^{i})=\bigcap_{i\in[m]}\operatorname{\textnormal{minx}}(t_{2}^{i})=\bigcap_{a\in\textnormal{pr}_{1}(S)}\operatorname{\textnormal{minx}}(a).$
Lemma 5.12 applied to the weakly connected component of $E$ containing $\textnormal{pr}_{1}(S)$ yields $\bigcap_{a\in\textnormal{pr}_{1}(S)}\operatorname{\textnormal{minx}}(a)\neq\varnothing.$ Therefore, the tuple $s\coloneqq{\operatorname{\textnormal{lex}}^{[m]}}(t^{1},\dots,t^{m})$ satisfies $\min(s_{1})=\min(s_{2})={\operatorname{\textnormal{lex}}^{[m]}}(c,\dots,c)$, as well as $\operatorname{\textnormal{minx}}(s_{1})=\bigcap_{a\in\textnormal{pr}_{1}(S)}\operatorname{\textnormal{minx}}(a)=\operatorname{\textnormal{minx}}(s_{2})$. Moreover, by the choice of $t^{1},\dots,t^{m}$ and injectivity of $\operatorname{\textnormal{lex}}^{[m]}$}, we also have $s\in S^{\prime}$. ∎
Lemma 5.14.
Let $R$ be an $n$-ary 22-transitive relation on $\mathbb{Q}^{k}$ that is preserved by $\operatorname{\textnormal{lex}}$. Either there is $t\in R$ with $|M(t)|=1,$ or $|M(t)|=n$ for all $t\in R$.
Proof.
The proof is similar to the one of Lemma 5.6, using $\operatorname{\textnormal{lex}}^{[n-1]}$} in place of $\operatorname{mi}^{[n-1]}$}. ∎
Lemma 5.15.
Let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ with $n\geq 3$. If $R$ is cyclic, $2$-transitive, and preserved by $\operatorname{\textnormal{lex}}$, then $R^{\prime}$ contains a min-clean tuple.
Proof.
If $R$ contains a tuple $t$ with $|M(t)|=1$, this tuple is min-clean. It then follows that for any $s\in R^{\prime}$ (which exists by Lemma 3.7), the tuple $\operatorname{\textnormal{lex}}(t,s)\in R^{\prime}$ is also min-clean. If $R$ does not contain such a tuple, then by Lemma 5.14 all tuples satisfy $|M(t)|=n$ and are, in particular, min-ready. Take $t\in R^{\prime}$ arbitrarily. As in the proof of Lemma 5.7, employing Lemma 5.11 instead of Lemma 5.3, one shows that $\bigcap_{i\leq n}\operatorname{\textnormal{minx}}(t_{i})\neq\varnothing.$ Let $t^{1},\dots,t^{n}$ denote all cyclic permutations of the tuple $t$. We have $t^{i}\in R$ for all $i\leq n$ by cyclicity of $R$. By injectivity of $\operatorname{\textnormal{lex}}^{[n]}$}, the tuple $s\coloneqq{\operatorname{\textnormal{lex}}^{[n]}}(t^{1},\dots,t^{n})$ is contained in $R^{\prime}$. Moreover, observe that $\min(s)={\operatorname{\textnormal{lex}}^{[m]}}(\min(t),\dots,\min(t))$ and $\operatorname{\textnormal{minx}}(s_{i})=\bigcap_{j\leq n}\operatorname{\textnormal{minx}}(t_{j})$ for all $i\leq n$. In particular, $s\in R^{\prime}$ is min-clean. ∎

### v2

@@
 1. Introduction
-1.1. Constraint Satisfaction Problems
-The Constraint Satisfaction Problem induced by a relational structure $\mathbb{A}$, denoted by $\operatorname{\textnormal{CSP}}(\mathbb{A})$, is the computational problem of deciding, given a finite input structure $\mathbb{B}$ of the same signature as $\mathbb{A}$, whether there exists a homomorphism from $\mathbb{B}$ to $\mathbb{A}$, i.e.
+The Constraint Satisfaction Problem induced by a relational structure $\mathbb{A}$, denoted by $\operatorname{CSP}(\mathbb{A})$, is the computational problem of deciding, given a finite input structure $\mathbb{B}$ of the same signature as $\mathbb{A}$, whether there exists a homomorphism from $\mathbb{B}$ to $\mathbb{A}$, i.e.
 a map preserving all relations.
@@
 This notion of fixed-template CSPs provides a uniform framework for modelling many classical computational problems such as graph-colouring problems, $3$-SAT, or solving equations.
-In 2017, two independent confirmations of a complexity dichotomy for CSPs induced by arbitrary finite structures – conjectured already in [FV93, FV98] – marked a breakthrough in the research programme on the complexity of finite-domain CSPs: for every finite structure $\mathbb{A}$, either $\operatorname{\textnormal{CSP}}(\mathbb{A})$ is solvable in polynomial time, or it is NP-complete [Bul17, Zhu17, Zhu20a], contrasting Ladner’s theorem [Lad75].
+In 2017, two independent confirmations of a complexity dichotomy for CSPs induced by arbitrary finite structures – conjectured already in [FV93, FV98] – marked a breakthrough in the research programme on the complexity of finite-domain CSPs: for every finite structure $\mathbb{A}$, either $\operatorname{CSP}(\mathbb{A})$ is solvable in polynomial time, or it is NP-complete [Bul17, Zhu17, Zhu20a], contrasting Ladner’s theorem [Lad75].
 Yet, prominent computational problems that can be phrased as the CSPs of a fixed template $\mathbb{A}$ require the domain of $\mathbb{A}$ to be of countably infinite size – for example, the problem of deciding whether a given finite digraph contains a cycle can be phrased as CSP over $(\mathbb{Q};<)$, but cannot be modelled in this way by any finite template.
@@
 1.1. The algebraic approach
-The complexity of finite-domain CSPs is determined by the ‘expressive power’ of the underlying structure $\mathbb{A}$, which is encoded in the set $\operatorname{\textnormal{Pol}}(\mathbb{A})$ of all compatible finitary operations – the polymorphism clone of $\mathbb{A}$.
+The complexity of finite-domain CSPs is determined by the ‘expressive power’ of the underlying structure $\mathbb{A}$, which is encoded in the set $\operatorname{Pol}(\mathbb{A})$ of all compatible finitary operations – the polymorphism clone of $\mathbb{A}$.
 The study of the computational complexity of CSPs via the polymorphisms they have is what is now understood as the algebraic approach to CSPs, first developed in [Jea98, BJK05] and unified by [BOP18].
-The finite-domain CSP dichotomy now takes the following form: as the only source of NP-completeness for $\operatorname{\textnormal{CSP}}(\mathbb{A})$ stands the ability of $\mathbb{A}$ to pp-construct EVERYTHING, i.e.
+The finite-domain CSP dichotomy now takes the following form: as the only source of NP-completeness for $\operatorname{CSP}(\mathbb{A})$ stands the ability of $\mathbb{A}$ to pp-construct EVERYTHING, i.e.
 every finite structure (see Section 2.1 for the definition of a pp-construction).
@@
 Every finite structure that lacks this property gives rise to a CSP solvable in polynomial time, and its polymorphism clone satisfies certain symmetries.
-Here, we say that the polymorphism clone $\operatorname{\textnormal{Pol}}(\mathbb{A})$ of a structure $\mathbb{A}$ satisfies an identity, if the identity is witnessed by operations contained in $\operatorname{\textnormal{Pol}}(\mathbb{A})$ for all evaluations of their arguments.
-For example, $\operatorname{\textnormal{Pol}}(\mathbb{A})$ always contains
+Here, we say that the polymorphism clone $\operatorname{Pol}(\mathbb{A})$ of a structure $\mathbb{A}$ satisfies an identity, if the identity is witnessed by operations contained in $\operatorname{Pol}(\mathbb{A})$ for all evaluations of their arguments.
+For example, $\operatorname{Pol}(\mathbb{A})$ always contains
   • a $6$-ary polymorphism $s$ witnessing the $6$-ary Siggers identity [Sig10]
@@
 An algebraic approach to CSPs via polymorphisms is for countably infinite structures only possible in the setting of $\omega$-categoricity.
-Indeed, if $\mathbb{A}$ is $\omega$-categorical, then the complexity of $\operatorname{\textnormal{CSP}}(\mathbb{A})$ is again captured within the ‘local’ algebraic structure of $\operatorname{\textnormal{Pol}}(\mathbb{A})$ [BP15b].
-Moreover, for first-order reducts of finitely bounded homogeneous structures $\mathbb{A}$, it is conjectured that omni-expressivity remains the only source of NP-completeness of $\operatorname{\textnormal{CSP}}(\mathbb{A})$ [BPP21, BOP18, BKO+17, BKO+19].
+Indeed, if $\mathbb{A}$ is $\omega$-categorical, then the complexity of $\operatorname{CSP}(\mathbb{A})$ is again captured within the ‘local’ algebraic structure of $\operatorname{Pol}(\mathbb{A})$ [BP15b].
+Moreover, for first-order reducts of finitely bounded homogeneous structures $\mathbb{A}$, it is conjectured that omni-expressivity remains the only source of NP-completeness of $\operatorname{CSP}(\mathbb{A})$ [BPP21, BOP18, BKO+17, BKO+19].
 Naturally, the question arises whether the identities (1)-(5) have counterparts for $\omega$-categorical structures that are not omni-expressive.
@@
 In the early stages of the systematic research programme on CSPs, those induced by finite graphs were among the first to be studied.
-Observe that the classical $k$-colouring problem coincides with $\operatorname{\textnormal{CSP}}(\mathbb{K}_{k})$, where $\mathbb{K}_{k}$ denotes the clique on $k$ vertices.
+Observe that the classical $k$-colouring problem coincides with $\operatorname{CSP}(\mathbb{K}_{k})$, where $\mathbb{K}_{k}$ denotes the clique on $k$ vertices.
 Vastly generalising the NP-completeness of the $k$-colouring problem for all $k\geq 3$, Hell and Nešetřil showed that every undirected non-bipartite graph is either omni-expressive (and hence its CSP is NP-complete), or has a loop (in which case it is entirely inexpressive, and in particular, its CSP trivial) [HN90].
@@
 Let $\mathbb{A}$ be a finite relational structure that is not omni-expressive.
-If $\mathbb{G}$ is any finite smooth digraph of algebraic length $1$, then $\operatorname{\textnormal{Pol}}(\mathbb{A})$ satisfies the $\mathbb{G}$-loop condition.
+If $\mathbb{G}$ is any finite smooth digraph of algebraic length $1$, then $\operatorname{Pol}(\mathbb{A})$ satisfies the $\mathbb{G}$-loop condition.
+1.3. Our contribution
 An oligomorphic subgroup $\Omega$ of the automorphism group of a countable digraph $\mathbb{G}$ has, in particular, finitely many orbits in its action on the domain of $\mathbb{G}$.
@@
 Observe that a loop in this finite digraph comes from an edge in $\mathbb{G}$ between two vertices belonging to the same $\Omega$-orbit – a so-called pseudo-loop modulo $\Omega$.
-The non-existence of pseudo-loops in $\mathbb{G}$ has been identified as a source of computational hardness for $\operatorname{\textnormal{CSP}}(\mathbb{G})$ in several cases [BP20, BBK+23, BKNP25].
+The non-existence of pseudo-loops in $\mathbb{G}$ has been identified as a source of computational hardness for $\operatorname{CSP}(\mathbb{G})$ in several cases [BP20, BBK+23, BKNP25].
 Following this direction, we prove a variant of Theorem 1.1 for the case of digraphs $\mathbb{G}$ that are pp-interpretable in a temporal constraint language $\mathbb{A}$ that is not omni-expressive.
@@
 The latter naturally acts on $\mathbb{A}$ by automorphisms, and in fact also on $\mathbb{G}$ by the definition of a (pp-)interpretation.
-Hence, we may regard $\operatorname{\textnormal{Aut}}((\mathbb{Q};<))$ also as an oligomorphic subgroup of $\operatorname{\textnormal{Aut}}(\mathbb{G})$.
+Hence, we may regard $\operatorname{Aut}((\mathbb{Q};<))$ also as an oligomorphic subgroup of $\operatorname{Aut}(\mathbb{G})$.
 Theorem 1.3.
 Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive.
-If $\mathbb{G}$ is any smooth digraph that is pp-interpretable in $\mathbb{A}$ and has pseudo-algebraic length $1$ modulo $\operatorname{\textnormal{Aut}}((\mathbb{Q};<))$, then $\mathbb{G}$ contains a pseudo-loop modulo $\operatorname{\textnormal{Aut}}((\mathbb{Q};<))$.
+If $\mathbb{G}$ is any smooth digraph that is pp-interpretable in $\mathbb{A}$ and has pseudo-algebraic length $1$ modulo $\operatorname{Aut}((\mathbb{Q};<))$, then $\mathbb{G}$ contains a pseudo-loop modulo $\operatorname{Aut}((\mathbb{Q};<))$.
 Theorem 1.3 reveals the dichotomous nature of temporal constraint languages in terms of their expressivity.
@@
 Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive.
-If $\mathbb{G}$ is any finite smooth digraph of algebraic length $1$, then $\operatorname{\textnormal{Pol}}(\mathbb{A})$ satisfies the $\mathbb{G}$-pseudo-loop condition.
+If $\mathbb{G}$ is any finite smooth digraph of algebraic length $1$, then $\operatorname{Pol}(\mathbb{A})$ satisfies the $\mathbb{G}$-pseudo-loop condition.
 Our result confirms that the full range of loop conditions whose satisfaction is known to characterise non-omni-expressivity in the finite setting lifts, in their pseudo-versions, to all not omni-expressive temporal structures.
@@
 As it turns out, sets of identities of the form as (3) and (4) are better suited for deriving further structural properties that may, in particular, be used in algorithms [BK09, BK14, Bul06, Zhu17, Zhu20a, MP22, MP24, KTV24, PRSS25].
-They are formalised through the notion of a $\mathbb{T}$-loop condition for hypergraphs $\mathbb{T}$ of arity $n\geq 3$, defined as a set of n1n-1 identities, explained in detail e.g.
+They are formalised through the notion of a $\mathbb{T}$-loop condition for hypergraphs $\mathbb{T}$ of arity $n\geq 3$, defined as a set of n1n-1 identities, as explained in detail e.g.
 in [GJP19].
-The following loop lemma for finite symmetric hypergraphs of suitable arity implies – in a similar way to how 1.2 is derived from Theorem 1.1 – the existence of WNU polymorphisms in all not omni-expressive finite structures.
-Here, a loop in an $n$-ary relation $R$ is a constant tuple $(a.\dots,a)\in R$.
-Theorem 1.5 (follows from [BK12], see also [Zhu20b]).
-Let $\mathbb{T}$ be a finite symmetric hypergraph of arity $n$ such that $p\nmid n$ for all prime numbers $p\leq|T|$. Either $\mathbb{T}$ is omni-expressive, or $\mathbb{T}$ contains a loop.
-Extending the approach used in the proof of Theorem 1.3 to higher arities, we proof a version of Theorem 1.5 for the case symmetric hypergraphs pp-interpretable in a not omni-expressive temporal template.
-Note that in contrast to the finite version, no number-theoretic restrictions on the arity of $\mathbb{T}$ apply.
+Extending our approach to higher arities, we prove a version of Theorem 1.3 for the case of hypergraphs pp-interpretable in a not omni-expressive temporal template.
+A pseudo-loop in an $n$-ary hypergraph is a hyperedge all of whose coordinates belong to the same orbit; a hypergraph is cyclic if its hyperedges are invariant under cyclic permutations, and it is $2$-transitive if it is invariant under the action of some $2$-transitive subgroup of the symmetric group on $n$ elements.
 Theorem 1.5.
 Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive.
-If $\mathbb{T}$ is any symmetric hypergraph of arity $n\geq 3$ that is pp-interpretable in $\mathbb{A}$, then $\mathbb{T}$ contains a pseudo-loop modulo $\operatorname{\textnormal{Aut}}((\mathbb{Q};<))$.
+If $\mathbb{T}$ is any cyclic and $2$-transitive hypergraph of arity $n\geq 3$ that is pp-interpretable in $\mathbb{A}$, then $\mathbb{T}$ contains a pseudo-loop modulo $\operatorname{Aut}((\mathbb{Q};<))$.
 The pseudo-loop condition induced by a hypergraph is again obtained by composing all identities of the corresponding loop-condition with unary function symbols.
-Since the hypergraphs inducing the Olšák- and the WNU-identities, respectively, are symmetric, the existence of pseudo-Olšák polymorphisms and pseudo-WNU polymorphisms of all arities $n\geq 3$ follows from Theorem 1.6 in the standard way.
+Since the hyperedges of the hypergraphs inducing the Olšák- and the WNU-identities are invariant under all permutations, they satisfy in particular the imposed symmetry conditions; hence, the existence of pseudo-Olšák polymorphisms and pseudo-WNU polymorphisms of all arities $n\geq 3$ follows from Theorem 1.5 in the standard way.
 Corollary 1.6.
 Let $\mathbb{A}$ be a temporal constraint language that is not omni-expressive.
-If $\mathbb{T}$ is any finite symmetric hypergraph of arity $n\geq 3$, then $\operatorname{\textnormal{Pol}}(\mathbb{A})$ satisfies the $\mathbb{T}$-pseudo-loop condition.
+If $\mathbb{T}$ is any finite cyclic and $2$-transitive hypergraph of arity $n\geq 3$, then $\operatorname{Pol}(\mathbb{A})$ satisfies the $\mathbb{T}$-pseudo-loop condition.
+1.4. Is $2$-transitivity the new cyclicity?
+Let us remark that the conclusion of Theorem 1.5 clearly fails in the case $n=2$, as it would amount to the existence of a binary polymorphism that is pseudo-commutative – yet temporal constraint languages preserved by the binary operations $\ell\ell$ or $\operatorname{mi}$ are neither omni-expressive, nor must they admit such polymorphisms.
+It might be of interest to point out that pseudo-loop lemmata involving symmetry conditions provided by non-trivial group actions have also been established for finite hypergraphs of arity $n\geq 3$ (see [BK12, Zhu20b, Bru23, Zhu24]).
+Most influentially, the study of finite cyclic hypergraphs in [BK12] stands as a landmark result in this context.
+In the temporal setting, cyclicity by itself is not a sufficiently strong condition: as mentioned before, there exist not-omni expressive temporal constraint languages that do not admit any pseudo-cyclic polymorphisms.
+A concrete example is the structure $(\mathbb{Q};<,x\neq y\vee u=v)$ [Bod21, Proposition], highlighting the necessity of the $2$-transitivity condition on top of the cyclicity condition in Theorem 1.5.
+Observe that these symmetry conditions are proper in the sense that a hypergraph may be cyclic and $2$-transitive without being fully symmetric, i.e.
+invariant under all permutations of its hyperedges.
+For example, for odd $n\geq 5$ the alternating group $A_{n}$ is a proper 2-transitive subgroup of the full symmetric group $S_{n}$ which contains all cyclic permutations.
+Whether or not cyclicity is truly a necessary condition in Theorem 1.5 remains an interesting open question for future work.
+For hypergraphs of arity $n=3$, it is readily implied by 2-transitivity, whereas for all arities $n>3$, it is not.
 2. Preliminaries
@@
 For $n\in\mathbb{N}$, we denote the set $\{1,\dots,n\}$ by $[n]$.
-For an $n$-tuple $a=(a_{1},\dots,a_{n})$, we write $a_{i}$ or $\textnormal{pr}_{i}(a)$ for its $i$-th coordinate.
+For an $n$-tuple $a=(a_{1},\dots,a_{n})$, we write $a_{i}$ or $\operatorname{pr}_{i}(a)$ for its $i$-th coordinate.
 By $\ker(a)$ we denote the subset of $n]\times[n]$ consisting of all pairs $(i,j)$ for which $a_{i}=a_{j}$.
 A relation $R$ on a set $A$ is a subset $R\subseteq A^{n}$ for some $n\in\mathbb{N}$, which is referred to as the arity of $R$.
-For a subset $I$ of $[n]$, we denote by $\textnormal{pr}_{I}(R)$ the $|I|$-ary relation obtained by projecting $R$ to all of its coordinates $i$ with $i\in I$.
-We also write $R|_{B_{1}\times\dots\times B_{n}}$ for the relation $R\cap(B_{1}\times\dots\times B_{n})$, where $B_{i}\subseteq A$.
+For a subset $I$ of $[n]$, we denote by $\operatorname{pr}_{I}(R)$ the $|I|$-ary relation obtained by projecting $R$ to all of its coordinates $i$ with $i\in I$.
 For the purpose of this paper, we assume all relations considered to be non-empty.
@@
 An automorphism of $\mathbb{A}$ is a bijective homomorphism $f:A\to A$ whose inverse mapping $f^{-1}$ is also a homomorphism.
-By $\operatorname{\textnormal{Aut}}(\mathbb{A})$ we denote the automorphism group of $\mathbb{A}$.
-In the following, by componentwise application we will understand all mappings also as functions on $n$-tuples.
+By $\operatorname{Aut}(\mathbb{A})$ we denote the automorphism group of $\mathbb{A}$.
+In the following, by componentwise application we will understand all mappings also as functions [Zhu24] on $n$-tuples.
 A first-order formula $\phi$ is called primitive positive (pp) over a set consisting of relational symbols $R_{1},\dots,R_{n}$ if it involves only the predicates $R_{i}$, existential quantification, and conjunction.
@@
 We say that $\mathbb{A}$ $p$-interprets $\mathbb{B}$ if there exist $k\geq 1$ and a partial surjective map $h:A^{k}\to B$ with the property that for every relation $R\subseteq B^{n}$ that is either a relation of $\mathbb{B}$, the equality relation on $B$, or $B$, its preimage $h^{-1}(R)$ – regarded as a relation of arity $nk$ on $A$ – is pp-definable in $\mathbb{A}$.
-If $\mathbb{A}$ $p$-interprets $\mathbb{B}$, then $\operatorname{\textnormal{CSP}}(\mathbb{B})$ is log-space reducible to $\operatorname{\textnormal{CSP}}(\mathbb{A})$ [BJK05].
+If $\mathbb{A}$ $p$-interprets $\mathbb{B}$, then $\operatorname{CSP}(\mathbb{B})$ is log-space reducible to $\operatorname{CSP}(\mathbb{A})$ [BJK05].
 We say that $\mathbb{A}$ $p$-constructs $\mathbb{B}$ if $\mathbb{B}$ is homomorphically equivalent to a structure that is pp-interpretable in $\mathbb{A}$.
-As homomorphically equivalent structures have the same CSPs, this also implies a log-space reduction from $\operatorname{\textnormal{CSP}}(\mathbb{B})$ to $\operatorname{\textnormal{CSP}}(\mathbb{A})$.
+As homomorphically equivalent structures have the same CSPs, this also implies a log-space reduction from $\operatorname{CSP}(\mathbb{B})$ to $\operatorname{CSP}(\mathbb{A})$.
 We say that $\mathbb{A}$ pp-constructs EVERYTHING if $\mathbb{A}$ pp-constructs every finite structure; we then call $\mathbb{A}$ omni-expressive.
@@
 By componentwise evaluation, $\Omega$ acts on $A^{k}$ for every $k\in\mathbb{N}.$ The group is oligomorphic if this action has finitely many orbits for every $k\in\mathbb{N}$.
-A countable structure $\mathbb{A}$ is $\omega$-categorical if $\operatorname{\textnormal{Aut}}(\mathbb{A})$ is oligomorphic.
+A countable structure $\mathbb{A}$ is $\omega$-categorical if $\operatorname{Aut}(\mathbb{A})$ is oligomorphic.
 2.2. Graph-theoretic
 A digraph is a relational structure of the form $\mathbb{G}=(G;E)$, where $E$ is binary.
-$\mathbb{G}$ is called smooth if $\textnormal{pr}_{1}(E)=\textnormal{pr}_{2}(E)$.
+$\mathbb{G}$ is called smooth if $\operatorname{pr}_{1}(E)=\operatorname{pr}_{2}(E)$.
+In this case, we also write $\operatorname{supp}(\mathbb{G})$ for the set $\operatorname{pr}_{1}(E)$.
 If $E$ is symmetric, $\mathbb{G}$ is also called a graph.
@@
 A digraph $\mathbb{G}$ is weakly connected if for all distinct $a,b\in G$ there exists an $E$-walk from $a$ to $b$.
-A subset $W\subseteq G$ is called a weakly connected component of $\mathbb{G}$ if it is maximal with respect to inclusion such that the digraph $\mathbb{G}|_{W}\coloneqq(W;E|_{W\times W})$ is weakly connected.
 A fence in $E$ from $a_{0}$ to $a_{2n}$ is an $E$-walk of the form $(a_{0}\dots a_{2n})$ where $(a_{2i-2},a_{2i-1})\in E$ and $(a_{2i-1},a_{2i})\in E^{-1}$ for all $i\leq n$.
@@
 For $m\in\mathbb{N}$, a fence in $E^{m}$ is also called an $m$-fence in $E$.
-We say that $\mathbb{G}$ is linked if there exists $m\in\mathbb{N}$ such that for any $a,b\in\textnormal{pr}_{1}(E)$ there exists an $m$-fence in $E$ from $a$ to $b$.
+We say that $\mathbb{G}$ is linked if there exists $m\in\mathbb{N}$ such that for any $a,b\in\operatorname{pr}_{1}(E)$ there exists an $m$-fence in $E$ from $a$ to $b$.
 A finite smooth digraph is linked if and only if it is weakly connected and has algebraic length $1$ [BK12, Claim 3.8].
 A hypergraph is any structure $\mathbb{T}=(T;R)$ containing only one relation $R$, where $R$ is not necessarily binary.
+We call elements of $R$ hyperedges.
 Let $n$ be the arity of $R$, and let $\Sigma$ be a subgroup of the symmetric group $\textnormal{Sym}(n)$ on $n$ elements.
 We say that $\mathbb{T}$ is $\Sigma$-invariant if for every $a\in R$ and $\pi\in\Sigma$ also $(a_{\pi(1)},\dots,a_{\pi(n)})\in R$.
-A hypergraph is symmetric if it is $\textnormal{Sym}(n)$-invariant, cyclic if it is invariant with respect to the group of cyclic shifts, and $2$-transitive if it is $\Sigma$-invariant for some $2$-transitive group $\Sigma$.
+A hypergraph is cyclic if it is invariant with respect to the group of cyclic shifts, and $2$-transitive if it is $\Sigma$-invariant for some $2$-transitive group $\Sigma$.
 A loop is a constant tuple $(a,\dots,a)\in R$.
@@
 For an $\omega$-categorical structure $\mathbb{A}$, a relation $R$ is pp-definable in $\mathbb{A}$ if and only if it is preserved by all polymorphisms of $\mathbb{A}$ [BN06].
-The polymorphism clone of $\mathbb{A}$ is the set of all polymorphism of $\mathbb{A}$, and it is denoted by $\operatorname{\textnormal{Pol}}(\mathbb{A})$.
+The polymorphism clone of $\mathbb{A}$ is the set of all polymorphism of $\mathbb{A}$, and it is denoted by $\operatorname{Pol}(\mathbb{A})$.
 Observe that it indeed forms a clone in the sense of universal algebra, as it is closed under composition and contains all projections.
@@
 A temporal relation is a relation first-order definable in $(\mathbb{Q};<)$.
-A temporal constraint language is a relational structure all of whose relations are temporal.
-Clearly, the natural action of $\operatorname{\textnormal{Aut}}(\mathbb{Q})\coloneqq\operatorname{\textnormal{Aut}}((\mathbb{Q};<))$ on $\mathbb{Q}^{k}$ has finitely many orbits for every $k\in\mathbb{N}$.
+A temporal constraint language is a relational structure with domain $\mathbb{Q}$ all of whose relations are temporal.
+Clearly, the natural action of $\operatorname{Aut}(\mathbb{Q})\coloneqq\operatorname{Aut}((\mathbb{Q};<))$ on $\mathbb{Q}^{k}$ has finitely many orbits for every $k\in\mathbb{N}$.
 The automorphisms of a structure preserve all relations first-order definable in it.
@@
 Let $F$ be a set of operations on $\mathbb{Q}$.
-We say that the clone generated by $F$ is the smallest clone of operations $\mathcal{C}$ that contains $F\cup\operatorname{\textnormal{Aut}}(\mathbb{Q})$ and is closed under interpolation; that is, an operation $g$ belongs to $\mathcal{C}$ if and only if for every finite subset $A\subseteq\mathbb{Q}$, there is $f\in\mathcal{C}$ agreeing with $g$ on $A$.
+We say that the clone generated by $F$ is the smallest clone of operations $\mathcal{C}$ that contains $F\cup\operatorname{Aut}(\mathbb{Q})$ and is closed under interpolation; that is, an operation $g$ belongs to $\mathcal{C}$ if and only if for every finite subset $A\subseteq\mathbb{Q}$, there is $f\in\mathcal{C}$ agreeing with $g$ on $A$.
 In the case that $F$ contains only one operation $f$, we say that $f$ generates $g$.
@@
 In this construction, one can easily enforce both that any of the endomorphisms has finitely many specified fixed points, or none at all.
+All operations that have the listed properties of $\operatorname{mi}$ generate the same clone, and the same holds for $\operatorname{mx}$.
 Note that our definition of $\operatorname{mi}$ differs slightly from the original one in [BK10].
-The two operations generate the same clone, and in particular each other.
+Operations $\operatorname{mi}$ defined either way generate the same clone, and in particular each other.
 Our definition of $\operatorname{mi}$ is in line with [BR22].
@@
 Let $\mathbb{A}$ be a temporal constraint language.
-If $\mathbb{A}$ is not omni-expressive, then it is preserved by one of $\min$, $\operatorname{mi}$, $\operatorname{mx}$, $\ell\ell$, their duals, or a constant operation, and $\operatorname{\textnormal{CSP}}(\mathbb{A})$ is solvable in polynomial time.
+If $\mathbb{A}$ is not omni-expressive, then it is preserved by one of $\min$, $\operatorname{mi}$, $\operatorname{mx}$, $\ell\ell$, their duals, or a constant operation, and $\operatorname{CSP}(\mathbb{A})$ is solvable in polynomial time.
 Figure 1. Polymorphisms of temporal constraint languages
 […]
-For the proofs of Theorems 1.3 and 1.6, we will additionally draw on two binary operations $\operatorname{\textnormal{lex}}$ and $pp$.
-As $\ell\ell$ and $\operatorname{dual}(\ell\ell)$ generate $\operatorname{\textnormal{lex}}$, and each of $\min,\operatorname{mi},$ and $\operatorname{mx}$ generates $pp$, it follows that one of $\operatorname{\textnormal{lex}}$, $p$,$p$, and $\operatorname{dual}(pp)$ is contained in the polymorphism clone of every not omni-expressive constraint language.
-It is worth noting, however, that the presence of $\operatorname{\textnormal{lex}}$ or $pp$ in a polymorphism clone does not, on its own, imply polynomial-time solvability of the underlying template.
-For example, the temporal constraint language modelling the classical Betweenness problem is preserved by $\operatorname{\textnormal{lex}}$, yet is NP-complete [Opa79].
+For the proofs of Theorems 1.3 and 1.5, we will additionally draw on two binary operations $\operatorname{lex}$ and $pp$.
+As $\ell\ell$ and $\operatorname{dual}(\ell\ell)$ generate $\operatorname{lex}$, and each of $\min,\operatorname{mi},$ and $\operatorname{mx}$ generates $pp$, it follows that one of $\operatorname{lex}$, $p$,$p$, and $\operatorname{dual}(pp)$ is contained in the polymorphism clone of every not omni-expressive constraint language.
+It is worth noting, however, that the presence of $\operatorname{lex}$ or $pp$ in a polymorphism clone does not, on its own, imply polynomial-time solvability of the underlying template.
+For example, the temporal constraint language modelling the classical Betweenness problem is preserved by $\operatorname{lex}$, yet is NP-complete [Opa79].
 Definition 2.4.
-Let $\operatorname{\textnormal{lex}}:\mathbb{Q}^{2}\to\mathbb{Q}$ be a binary operation satisfying $\operatorname{\textnormal{lex}}(x,y)\leq\operatorname{\textnormal{lex}}(x^{\prime},y^{\prime})$ if and only if
+Let $\operatorname{lex}:\mathbb{Q}^{2}\to\mathbb{Q}$ be a binary operation satisfying $\operatorname{lex}(x,y)\leq\operatorname{lex}(x^{\prime},y^{\prime})$ if and only if
   • $x<x^{\prime}$ or
@@
 It is easy to see that all operations with these properties generate the same clone [BK10, Observation 1].
-As $\ell\ell$ and $\operatorname{dual}(\ell\ell)$ also satisfy them when restricted to $\{(x,y)\in\mathbb{Q}\times\mathbb{Q}\colon x,y<0\}$ and $\{(x,y)\in\mathbb{Q}\times\mathbb{Q}\colon x,y>0\}$, respectively, both of them generate $\operatorname{\textnormal{lex}}$.
+As $\ell\ell$ and $\operatorname{dual}(\ell\ell)$ also satisfy them when restricted to $\{(x,y)\in\mathbb{Q}\times\mathbb{Q}\colon x,y<0\}$ and $\{(x,y)\in\mathbb{Q}\times\mathbb{Q}\colon x,y>0\}$, respectively, both of them generate $\operatorname{lex}$.
 Note that lex is, by definition, injective.
@@
 Again, $pp_{q}$ generates $pp_{p}$ for $p\neq q$. By $pp$, we denote the operation $pp_{0}.$ It is generated from $\min,\operatorname{mx},$ and $\operatorname{mi}$ [BK10, Lemma 23].
-For every $k\in\mathbb{N}$, we denote by $\sim_{k}$ the orbit-equivalence of $\operatorname{\textnormal{Aut}}(\mathbb{Q})\curvearrowright\mathbb{Q}^{k}$.
+For every $k\in\mathbb{N}$, we denote by $\sim_{k}$ the orbit-equivalence of $\operatorname{Aut}(\mathbb{Q})\curvearrowright\mathbb{Q}^{k}$.
 In other words, for $x,y\in\mathbb{Q}^{k}$ we have $x\sim_{k}y$ if and only if $x_{i}\leq x_{j}\leftrightarrow y_{i}\leq y_{j}$ for all $i,j\leq k$.
@@
 Finally, we note that if $\mathbb{G}$ is pp-interpretable in a temporal constraint language $\mathbb{A}$, then every polymorphism of $\mathbb{A}$ induces a polymorphism of $\mathbb{G}$ [BJK05]: this follows from the fact that we can regard $\mathbb{G}$ as a structure whose domain is a pp-definable subset of some power of $\mathbb{Q}$ factored by a pp-definable equivalence relation, with all its relations being pp-definable as well.
-In particular, this is true for the automorphisms of $\mathbb{Q}$, hence $\operatorname{\textnormal{Aut}}(\mathbb{Q})\curvearrowright\mathbb{G}$.
-If $\mathbb{G}=(G;E)$ is a digraph, a pseudo-loop in $\mathbb{G}$ modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ is an edge whose vertices are contained in the same orbit of $\operatorname{\textnormal{Aut}}(\mathbb{Q})\curvearrowright\mathbb{G}$.
-We say that $\mathbb{G}$ has pseudo-algebraic length $1$ modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ if there exists an $E$-walk of algebraic length $1$ so that its start point and its end point are contained in the same $\operatorname{\textnormal{Aut}}(\mathbb{Q})$-orbit.
-Similarly, a pseudo-loop modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ in an $n$-ary hypergraph $\mathbb{T}=(T;R)$ that is pp-interpretable in a temporal constraint language is an edge $(a_{1},\dots,a_{n})\in R$ such that all $a_{i}$ belong to the same $\operatorname{\textnormal{Aut}}(\mathbb{Q})$-orbit.
+In particular, this is true for the automorphisms of $\mathbb{Q}$, hence $\operatorname{Aut}(\mathbb{Q})\curvearrowright\mathbb{G}$.
+If $\mathbb{G}=(G;E)$ is a digraph, a pseudo-loop in $\mathbb{G}$ modulo $\operatorname{Aut}(\mathbb{Q})$ is an edge whose vertices are contained in the same orbit of $\operatorname{Aut}(\mathbb{Q})\curvearrowright\mathbb{G}$.
+We say that $\mathbb{G}$ has pseudo-algebraic length $1$ modulo $\operatorname{Aut}(\mathbb{Q})$ if there exists an $E$-walk of algebraic length $1$ so that its start point and its end point are contained in the same $\operatorname{Aut}(\mathbb{Q})$-orbit.
+Similarly, a pseudo-loop modulo $\operatorname{Aut}(\mathbb{Q})$ in an $n$-ary hypergraph $\mathbb{T}=(T;R)$ that is pp-interpretable in a temporal constraint language is an edge $(a_{1},\dots,a_{n})\in R$ such that all $a_{i}$ belong to the same $\operatorname{Aut}(\mathbb{Q})$-orbit.
 3. Chasing Orbits
-For the proofs of Theorems 1.3 and 1.6, a pseudo-loop is constructed via a recursive procedure that, at each step, aligns the minimal elements of appropriately chosen tuples.
-We only consider relations on some power of $\mathbb{Q}$ that are preserved by one of $\min,\operatorname{mi},\operatorname{mx},$ or $\ell\ell$.
-The general case follows easily as discussed in the proof of Theorem 1.3.
+For the proofs of Theorems 1.3 and 1.5, a pseudo-loop is constructed via a recursive procedure that, at each step, aligns the minimal elements of appropriately chosen tuples.
+We only consider relations on some power of $\mathbb{Q}$ that are preserved by one of the operations $\min,\operatorname{mi},\operatorname{mx},$ or $\ell\ell$.
+The general case for relations pp-interpretable in temporal constraint languages easily reduces to this setting, as discussed in the proof of Theorem 1.3.
+In what follows, we fix an integer $k\geq 2$.
 Definition 3.1.
-Let $k\geq 1$.
-  • For every $a\in\mathbb{Q}^{k}$, we denote by $\min(a)$ the minimal entry of $a$. By $\operatorname{\textnormal{minx}}(a)$ we denote the set of all indices $i\leq k$ for which $a_{i}=\min(a)$.
-  • For all $a,b\in\mathbb{Q}^{k}$ and $I\subseteq[k]$, we write $a\sim_{I}b$ if $\textnormal{pr}_{I}(a)\sim_{|I|}\textnormal{pr}_{I}(b)$.
+x
+  • For every $a\in\mathbb{Q}^{k}$, we denote by $\min(a)$ the minimal entry of $a$. By $\operatorname{minx}(a)$ we denote the set of all indices $i\leq k$ for which $a_{i}=\min(a)$.
+  • For all $a,b\in\mathbb{Q}^{k}$ and $I\subseteq[k]$, we write $a\sim_{I}b$ if $\operatorname{pr}_{I}(a)\sim_{|I|}\operatorname{pr}_{I}(b)$.
   • For $1\leq m\leq k$ and $a\in\mathbb{Q}^{k}$, we set $I_{m}(a)$ for the set of those indices $i$ such that $a_{i}$ is among the $m$ smallest values that appear in $a$.
@@
   3. (3)
-    $\textnormal{pr}_{I_{m}(a^{m+1}_{i})}(a^{m+1}_{i})\sim_{m}a^{m}_{i}$, i.e.
+    $\operatorname{pr}_{I_{m}(a^{m+1}_{i})}(a^{m+1}_{i})\sim_{m}\operatorname{pr}_{I_{m}(a^{m}_{i})}(a^{m}_{i})$, i.e.
 the $\sim_{m}$-class of the subtuple of $a^{m+1}_{i}$ containing the $m$ smallest entries of $a^{m+1}_{i}$ coincides with the $\sim_{m}$-class of the subtuple of $a^{m}_{i}$ containing all of its $m$ smallest entries.
-After at most $k$-many steps, this recursion yields a tuple all of whose coordinates lie in the same orbit of $\operatorname{\textnormal{Aut}}(\mathbb{Q})\curvearrowright\mathbb{Q}^{k}$, i.e.
-a pseudo-loop in $R$.
+After at most $k$-many steps, this recursion yields a tuple all of whose coordinates lie in the same orbit of $\operatorname{Aut}(\mathbb{Q})\curvearrowright\mathbb{Q}^{k}$, i.e. a pseudo-loop in $R$.
 Definition 3.2.
@@
 By $M(a)$ we denote the set consisting of all $i\in[n]$ for which $\min(a_{i})=\min(a)$.
-We say that $a$ is min-clean if for all $i,j\in M(a)$ we have $\operatorname{\textnormal{minx}}(a_{i})=\operatorname{\textnormal{minx}}(a_{j})$.
-The technical part of the proofs of Theorems 1.3 and 1.6 concerns the existence of min-clean tuples, and is deferred to Section 5.
+We say that $a$ is min-clean if for all $i,j\in M(a)$ we have $\operatorname{minx}(a_{i})=\operatorname{minx}(a_{j})$.
+The technical part of the proofs of Theorems 1.3 and 1.5 concerns the existence of min-clean tuples, and is deferred to Section 6.
 In the following, we construct a pseudo-loop under the assumption that min-clean tuples exist.
 We note an asymmetry between the cases distinguished by Theorem 2.3.
-When the template is preserved by $\ell\ell$, the existence of min-clean tuples only relies on the polymorphism $\operatorname{\textnormal{lex}}$ (which is generated by $\ell\ell$), while the subsequent construction uses the full strength of $\ell\ell$.
+When the template is preserved by $\ell\ell$, the existence of min-clean tuples only relies on the polymorphism $\operatorname{lex}$ (which is generated by $\ell\ell$), while the subsequent construction uses the full strength of $\ell\ell$.
 In the other cases, the construction proceeds using $pp$ (which alone does not prevent omni-expressivity), and the existence of min-clean tuples is shown by employing the polymorphisms $\min,\operatorname{mi}$, and $\operatorname{mx}$, respectively.
 The construction is made up of a nested composition of $\ell\ell$ and $pp$, building on the following easy observation: let $a\in\mathbb{Q}^{k}$, $q\in\mathbb{Q}$, and define $J\coloneqq\{i\in[k]:a_{i}\leq q\}$.
-For every $b\in\mathbb{Q}^{k}$, the $\sim_{k}$-class of the $k$-tuple $pp_{q}(a,b)$ is determined only by the $\sim_{|J|}$-class of the $|J|$-tuple $\textnormal{pr}_{J}(a)$ and the $\sim_{(k-|J|)}$-class of the (k|J|)(k-|J|)-tuple $\textnormal{pr}_{[k]\setminus J}(b)$.
+For every $b\in\mathbb{Q}^{k}$, the $\sim_{k}$-class of the $k$-tuple $pp_{q}(a,b)$ is determined only by the $\sim_{|J|}$-class of the $|J|$-tuple $\operatorname{pr}_{J}(a)$ and the $\sim_{(k-|J|)}$-class of the (k|J|)(k-|J|)-tuple $\operatorname{pr}_{[k]\setminus J}(b)$.
 The same holds true for the tuple ${\ell\ell}_{q}(a,b)$, provided that $\ker(a)=\ker(b)$.
-For a binary operation $f$ and $m\geq 2$, we denote by $f^{[m]}$} the $m$-ary operation
-${f^{[m]}}(x_{1},\dots,x_{m})\coloneqq f(x_{1},f(x_{2},\dots,f(x_{m-1},x_{m})\dots)).$
-Note that $f^{[m]}$} is lies within every clone containing $f$.
+To simplify notation, for a binary operation $f$ and $m\geq 2$, we denote by $f^{[m]}$} the $m$-ary operation defined as ${f^{[m]}}(x_{1},\dots,x_{m})\coloneqq f(x_{1},f(x_{2},\dots,f(x_{m-1},x_{m})\dots)).$ Note that $f^{[m]}$} is lies within every clone containing $f$.
 3.1. Templates with pppp
 Lemma 3.3.
-Let $E$ be a binary smooth relation on $\mathbb{Q}^{k}$ that is preserved by $\langle pp\rangle$, and assume that $E$ contains a min-clean tuple $\binom{u^{1}}{v^{1}}$.
-Then there exist $I\subseteq k$, $m,n\geq 1,q\in\mathbb{Q}$, and $\binom{u^{2}}{v^{2}},\ldots,\binom{u^{n}}{v^{n}}\in E$ such that for all $u,v\in\mathbb{Q}^{k}$ the tuple defined by
+Let $E$ be a binary smooth relation on $\mathbb{Q}^{k}$ that is preserved by $\langle pp\rangle$.
+If $E$ contains a min-clean tuple $\binom{u^{1}}{v^{1}}$, then there exist $I\subseteq k$, $m,n\geq 1,q\in\mathbb{Q}$, and $\binom{u^{2}}{v^{2}},\ldots,\binom{u^{n}}{v^{n}}\in E$ such that for all $u,v\in\mathbb{Q}^{k}$ the tuple defined by
 $\binom{u^{\prime}}{v^{\prime}}\coloneqq{pp_{q}^{[n+1]}}\left(\binom{u^{1}}{v^{1}},\binom{u^{2}}{v^{2}},\dots,\binom{u^{n}}{v^{n}},\binom{u}{v}\right)$
@@
 Proof.
-If $\min u^{1}=\min v^{1}$ then $\operatorname{\textnormal{minx}}u^{1}=\operatorname{\textnormal{minx}}v^{1}$.
-In this case, we set $I\coloneqq\operatorname{\textnormal{minx}}u^{1}$, $m\coloneqq 1$, $n\coloneqq 1$, and $q\coloneqq\min u^{1}$.
+If $\min u^{1}=\min v^{1}$ then $\operatorname{minx}u^{1}=\operatorname{minx}v^{1}$.
+In this case, we set $I\coloneqq\operatorname{minx}u^{1}$, $m\coloneqq 1$, $n\coloneqq 1$, and $q\coloneqq\min u^{1}$.
 Otherwise, assume without loss of generality that $\min u^{1}<\min v^{1}$.
 We set $q\coloneqq\min u^{1}$, and inductively set $v^{i+1}\coloneqq u^{i}$, and pick $u^{i+1}$ arbitrarily so that $\binom{u^{i+1}}{v^{i+1}}\in T$.
-Set $J^{1}\coloneqq\operatorname{\textnormal{minx}}u^{1}$, and inductively $J^{i+1}\coloneqq\{j\;|\;u^{i+1}_{j}\leq q\}\setminus J^{1}\cup\cdots\cup J^{i}$.
+Set $J^{1}\coloneqq\operatorname{minx}u^{1}$, and inductively $J^{i+1}\coloneqq\{j\;|\;u^{i+1}_{j}\leq q\}\setminus J^{1}\cup\cdots\cup J^{i}$.
 Finally, we pick $n\geq 2$ minimal such that $J^{n}$ is empty, set $I$ to be the union $J^{1}\cup\cdots\cup J^{n}$, and $m\coloneqq|I|$.
 For all $i$ we have $u^{\prime}\sim_{J_{i}}u^{i}$ and $v^{\prime}\sim_{J_{i}}v^{i+1}=u^{i}$, hence $u^{\prime}\sim_{J_{i}}v^{\prime}$.
-Moreover $\textnormal{pr}_{J_{i}}(u^{\prime})<\textnormal{pr}_{J_{i+1}}(u^{\prime})$, and similarly $\textnormal{pr}_{J_{i}}(v^{\prime})<\textnormal{pr}_{J_{i+1}}(v^{\prime})$ for all $i$.
+Moreover $\operatorname{pr}_{J_{i}}(u^{\prime})<\operatorname{pr}_{J_{i+1}}(u^{\prime})$, and similarly $\operatorname{pr}_{J_{i}}(v^{\prime})<\operatorname{pr}_{J_{i+1}}(v^{\prime})$ for all $i$.
 It follows that $u^{\prime}\sim_{I}v^{\prime}$.
-Moreover, $\textnormal{pr}_{I}(u^{\prime})<\textnormal{pr}_{[k]\setminus I}(u^{\prime})$ and $\textnormal{pr}_{I}(v^{\prime})<\textnormal{pr}_{[k]\setminus I}(v^{\prime})$, hence $I_{m}(u^{\prime})=I_{m}(v^{\prime})=I$.
+Moreover, $\operatorname{pr}_{I}(u^{\prime})<\operatorname{pr}_{[k]\setminus I}(u^{\prime})$ and $\operatorname{pr}_{I}(v^{\prime})<\operatorname{pr}_{[k]\setminus I}(v^{\prime})$, hence $I_{m}(u^{\prime})=I_{m}(v^{\prime})=I$.
 Finally, it is easy to verify that $u^{\prime}\sim_{[k]\setminus I}u$ and $v^{\prime}\sim_{[k]\setminus I}v$.
 ∎
-A similar statement holds true for higher-arity relations $R$ that are symmetric. In fact, provided the existence of min-clean tuples in $R$, it suffices that $R$ is cyclic.
+A similar statement holds true for higher-arity relations $R$ that are cyclic:
 Lemma 3.4.
-Let $R$ be an $n$-ary cyclic relation on $\mathbb{Q}^{k}$ that is preserved by $\langle pp\rangle$, and assume that $R$ contains a min-clean tuple $u^{1}=(u^{1}_{1}\dots,u^{1}_{n})$.
-Then there exist $I\subseteq[k],q\in\mathbb{Q}$, and $u^{2},\dots,u^{n}\in R$ such that for all $u=(u_{1},\dots,u_{n})\in R$ the tuple $u^{\prime}=(u^{\prime}_{1},\dots,u^{\prime}_{n})$ defined by
+Let $n\geq 2$, and let $R$ be an $n$-ary cyclic relation on $\mathbb{Q}^{k}$ that is preserved by $\langle pp\rangle$.
+If $R$ contains a min-clean tuple $u^{1}$, then there exist $I\subseteq[k],q\in\mathbb{Q}$, and $u^{2},\dots,u^{n}\in R$ such that for all $u\in R$ the tuple $u^{\prime}$ defined by
 $u^{\prime}\coloneqq{pp_{q}^{[n+1]}}\left(u^{1},u^{2},\dots,u^{n},u\right)$
@@
 Without loss of generality assume that $1\in M(u^{1})$, and set $q\coloneqq\min(u^{1}_{1})$.
-By min-cleanliness of $u^{1}$, its cyclic permutations $u^{2}\coloneqq(u^{1}_{n},u^{1}_{1},u^{1}_{2},\dots),\dots,u^{n}\coloneqq(u^{1}_{2},\dots,u^{1}_{1})$ and $I\coloneqq\operatorname{\textnormal{minx}}(u^{1}_{1})$ now have the required properties.
+By min-cleanliness of $u^{1}$, its cyclic permutations $u^{2}\coloneqq(u^{1}_{2},u^{1}_{3},\dots,u^{1}_{1}),\dots,u^{n}\coloneqq(u^{1}_{n},u^{1}_{1},u^{1}_{2},\dots)$ and $I\coloneqq\operatorname{minx}(u^{1}_{1})$ now have the required properties.
 ∎
-We now get a pseudo-loop by inductively shrinking $k$.
-To ensure the existence of min-clean tuples in cyclic relations $R$ of arity $n\geq 3$, as shown in Section 5, we additionally require that $R$ be $\Sigma$-invariant for some $2$-transitive subgroup $\Sigma$ of $\textnormal{Sym}(n)$.
+In both cases, we now get a pseudo-loop by inductively shrinking $k$.
+To ensure the existence of min-clean tuples in cyclic relations $R$ of arity $n\geq 3$, as shown in Section 6, we need the additional requirement of 2-transitivity.
 Lemma 3.5.
@@
 Assume we have already shown it for all $j<k$.
-Lemmata 5.1, 5.5, 5.9, 5.2, 5.7, and 5.10 provide min-clean tuples in all cases, and the fact that $pp$ is contained in all three clones in question allows us to use Lemmata 3.3 and 3.4, respectively.
+Lemmata 6.6, 6.3, 6.9, 6.7, 6.5, and 6.10 provide min-clean tuples in all cases, and the fact that $pp$ is contained in all three clones in question allows us to use Lemmata 3.3 and 3.4, respectively.
 We prove the statement for the binary case; the case $n\geq 3$ follows by analogy.
 Let $\binom{u^{1}}{v^{1}}\in R$ be min-clean, and take $I,m,n,q,$ and $\binom{u^{2}}{v^{2}},\ldots,\binom{u^{n}}{v^{n}}\in R$ as in Lemma 3.3.
-By induction hypothesis applied to the relation $\textnormal{pr}_{[k]\setminus I}R$, there exists $\binom{u}{v}\in R$ such that $u\sim_{[k]\setminus I}v$.
+By induction hypothesis applied to the relation $\operatorname{pr}_{[k]\setminus I}(R)$, there exists $\binom{u}{v}\in R$ such that $u\sim_{[k]\setminus I}v$.
 For $\binom{u^{\prime}}{v^{\prime}}$ defined as in Lemma 3.3, we then have:
@@
 3.2. Templates with $\ell\ell$
-In order to adapt Lemma 3.3 to the setting of digraphs preserved by $\ell\ell$, we need to restrict to tuples that are ‘as injective as possible’.
+For the proof of Lemma 3.3, we make heavy use of the fact that $pp_{q}(x,y)$ depends solely on the value of $x$ when $x\leq q$, and on the value of $y$ when $x>q$.
+In contrast, every operation contained in the clone $\langle\ell\ell\rangle$ is injective in all its non-dummy variables.
+In order to adapt to this setting, we need to ‘synchronise’ the kernels of the tuples we work with.
 Definition 3.6.
-Let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$. For $S\subseteq R$, we set
-$S^{\prime}\coloneqq\{t\in\langle S\rangle_{\operatorname{\textnormal{lex}}}\;|\;\ker t=\bigcap_{s\in S}\ker s\}.$
-The relation $S^{\prime}$ inherits desirable properties from $S$:
+For an $n$-ary relation $R$ on $\mathbb{Q}^{k}$, set $R^{\prime}\coloneqq\{t\in\langle{R}\rangle_{\operatorname{lex}}\;|\;\ker t=\bigcap_{r\in R}\ker r\}.$
+The relation $R^{\prime}$ inherits desirable properties from $R$:
 Lemma 3.7.
-If $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{\textnormal{lex}}\rangle$.
-If $S\subseteq R$ is non-empty, then $S^{\prime}\neq\varnothing$.
-If $n=2$ and $S$ is smooth, then so is $S^{\prime}$.
-If $\Sigma$ is a subgroup of $\textnormal{Sym}(n)$ and $S$ is $\Sigma$-invariant, then so is $S^{\prime}$.
+Let $n\geq 2$, and let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$.
+Then $R^{\prime}\neq\varnothing$.
+If $n=2$ and $R$ is smooth, then so is $R^{\prime}$.
+If $\Sigma$ is a subgroup of $\textnormal{Sym}(n)$ and $R$ is $\Sigma$-invariant, then so is $R^{\prime}$.
 Proof.
-The first statement is obtained by applying $\operatorname{\textnormal{lex}}^{[m]}$} of a sufficiently high arity $m$ to tuples in $S$ until the kernel of the resulting tuple can no more be refined.
-Suppose that $n=2$ and $S$ is smooth.
-To see that $S^{\prime}$ is smooth, pick $t\in S^{\prime}$, $f\in\langle\operatorname{\textnormal{lex}}\rangle$ of some arity $m\in\mathbb{N}$, and $s^{1},\dots,s^{m}\in S$ such that $t=f(s^{1},\dots,s^{m}).$ We show that $t_{1}\in\textnormal{pr}_{2}(S^{\prime})$, the converse statement for $t_{2}$ is shown analogously.
-By smoothness of $S$, for every $i\leq m$ there exists $r^{i}\in\mathbb{Q}^{k}$ such that $\binom{r^{i}}{s^{i}_{1}}\in S$.
+The first statement is obtained by applying $\operatorname{lex}^{[m]}$} of a sufficiently high arity $m$ to tuples in $R$ until the kernel of the resulting tuple can no more be refined.
+Suppose that $n=2$ and $R$ is smooth.
+To see that $R^{\prime}$ is smooth, pick $t\in R^{\prime}$, $f\in\langle\operatorname{lex}\rangle$ of some arity $m\in\mathbb{N}$, and $s^{1},\dots,s^{m}\in R$ such that $t=f(s^{1},\dots,s^{m}).$ We show that $t_{1}\in\operatorname{pr}_{2}(R^{\prime})$, the converse statement for $t_{2}$ is shown analogously.
+By smoothness of $R$, for every $i\leq m$ there exists $r^{i}\in\mathbb{Q}^{k}$ such that $\binom{r^{i}}{s^{i}_{1}}\in R$.
 Let $w\coloneqq f(r^{1},\dots,r^{m})$.
-Picking any $\binom{u}{v}\in S^{\prime}$, we now have that $\operatorname{\textnormal{lex}}(\binom{w}{t_{1}},\binom{u}{v})\in S^{\prime}$.
-From $t_{1}\sim_{k}\operatorname{\textnormal{lex}}(t_{1},v)$, the statement follows.
-The last statement is clear from the definition of $S^{\prime}$.
+Picking any $\binom{u}{v}\in R^{\prime}$, we now have that $\operatorname{lex}(\binom{w}{t_{1}},\binom{u}{v})\in R^{\prime}$.
+From $t_{1}\sim_{k}\operatorname{lex}(t_{1},v)$, the statement follows.
+The last statement is clear from the definition of $R^{\prime}$.
 ∎
-The following is an adaptation of Lemma 3.3 to the setting of smooth binary relations preserved by $\langle\ell\ell\rangle$.
+For relations $R$ preserved by $\langle\ell\ell\rangle$, the existence of pseudo-loops is shown in a similar way as for relations preserved by $\langle pp\rangle$.
+Namely, it follows from adaptions of Lemmata 3.3 and 3.4 to this new setting (Lemmata 3.8 and 3.10 below) along with the existence of min-clean tuples (Lemmata 6.13 and 6.15, shown in Section 6).
+The crucial difference to the previous setting is that all tuples we work with must be contained in $R^{\prime}$.
+This additional prerequisite leads to the slightly weaker analogues to Lemmata 3.3 and 3.4, compensated by strengthened results showing the existence of min-clean tuples even within $R^{\prime}$.
 Lemma 3.8.
@@
 ∎
-Again, the existence of min-clean tuples as provided by Lemma 5.13 directly implies the existence of a pseudo-loop.
+In the proof of Theorem 1.3, we apply Lemma 3.8 to a smooth weakly connected relation $S$ of pseudo-algebraic length $1$ that is contained in the edge relation of a digraph preserved by $\langle\ell\ell\rangle$.
+We now conclude on the existence of a pseudo-loop within $\langle{S^{\prime}}\rangle_{\ell\ell}$.
 Lemma 3.9.
-Let $E$ be a binary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\ell\ell\rangle$.
-If $S\subseteq E$ is smooth, has pseudo-algebraic length $1$, and $\textnormal{pr}_{1}(S)$ is weakly connected, then $S^{\prime}$ contains a pseudo-loop.
+Let $S$ be a binary relation on $\mathbb{Q}^{k}$. If $S$ is smooth, weakly connected and of pseudo-algebraic length $1$, then $\langle{S^{\prime}}\rangle_{\ell\ell}$ contains a pseudo-loop.
 Proof.
 We repeat the induction in the proof of Lemma 3.5. This time, Lemma 6.13 provides a min-clean tuple within $S^{\prime}$, and we use Lemma 3.8 instead of Lemma 3.3 to finish the proof. ∎
-We proceed with the corresponding statement for hypergraphs preserved by $\langle\ell\ell\rangle$ that are $\Sigma$-invariant for some $2$-transitive cyclic group $\Sigma$.
+Just as for templates with $pp$, a statement similar to Lemma 3.8 also applies to higher-arity relations that are cyclic.
+Lemma 3.10.
+Let $n\geq 2$, and let $R$ be an $n$-ary cyclic relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\ell\ell\rangle$.
+If $R^{\prime}$ contains a min-clean tuple $u^{1}$, then there exist $I\subseteq[k],q\in\mathbb{Q}$, and $u^{2},\dots,u^{n}\in R$ such that for all $u\in R^{\prime}$ the tuple $u^{\prime}$ defined by
+$u^{\prime}\coloneqq{\ell\ell_{q}^{[n+1]}}\left(u^{1},u^{2},\dots,u^{n},u\right)$
+satisfies:
+  • $I_{1}(u^{\prime}_{i})=I$ for all $i\leq n$;
+  • $u^{\prime}_{i}\sim_{[k]\setminus I}u_{i}$ for all $i\leq n$.
+Proof.
+By Lemma 3.7, $R^{\prime}$ is cyclic, thus containing all cyclic permutations of $u^{1}$. The proof now is analogous to the one of Lemma 3.4. ∎
+Under the additional assumption of $2$-transitivity, we obtain a pseudo-loop also in this setting.
 Lemma 3.11.
@@
 Proof.
-Once more, we proceed by induction on $k$, employing Lemma 5.15 for the existence of a min-clean tuple within $R^{\prime}$, and the slight adaption of Lemma 3.4 to the setting of $\ell\ell$ changing $R$ to $R^{\prime}$.
-∎
-3.3. Proof of the main theorems
+We proceed by induction on $k$, employing Lemma 6.15 for the existence of a min-clean tuple within $R^{\prime}$ together with Lemma 3.10. ∎
+4. Proof of the main theorems
 With the classification of not omni-expressive temporal constraint languages in hand, the remaining task is to combine the results established thus far to prove our main theorems.
@@
 Replacing $\mathbb{G}$ by the preimage $\bar{\mathbb{G}}$ of $\mathbb{G}$ under the map of the pp-interpretation in $\mathbb{A}$, we may assume that $\mathbb{G}$ is a digraph defined on a finite power of $\mathbb{Q}$ whose edge relation $E$ is preserved by every polymorphism of $\mathbb{A}$.
-Indeed, by definition of a pp-interpretation, the preimage $\bar{E}$ of $E$ is pp-definable in $\mathbb{A}$, and therefore preserved by $\operatorname{\textnormal{Pol}}(\mathbb{A})$ [BN06].
-A pseudo-loop in $\bar{E}$ yields a pseudo-loop in $E$ modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$.
-Moreover, $\bar{E}$ is smooth and has pseudo-algebraic length modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$.
+Indeed, by definition of a pp-interpretation, the preimage $\bar{E}$ of $E$ is pp-definable in $\mathbb{A}$, and therefore preserved by $\operatorname{Pol}(\mathbb{A})$ [BN06].
+A pseudo-loop in $\bar{E}$ yields a pseudo-loop in $E$ modulo $\operatorname{Aut}(\mathbb{Q})$.
+Moreover, $\bar{E}$ is smooth and has pseudo-algebraic length modulo $\operatorname{Aut}(\mathbb{Q})$.
 By Theorem 2.3, $\mathbb{A}$ is preserved by one of $\ell\ell$, $\min$, $\operatorname{mi}$, $\operatorname{mx}$, their duals, or a constant operation.
@@
 Recalling that, for $f\in\{\ell\ell,\min,\operatorname{mi},\operatorname{mx}\}$, $\mathbb{G}$ is preserved by $\langle\operatorname{dual}(f)\rangle$ if and only if $\mathbb{G}$ is preserved by $\langle f\rangle$, this allows us to restrict ourselves to the cases where $\mathbb{G}$ is preserved by one of $\langle\ell\ell\rangle$, $\langle\min\rangle$, $\langle\operatorname{mi}\rangle$, and $\langle\operatorname{mx}\rangle$.
-In the first case, let $W$ be a weakly connected component of $\mathbb{G}$ of pseudo-algebraic length $1$ modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$, and set $S\coloneqq E|_{W\times W}$.
-By Lemma 3.9, $S^{\prime}\subseteq E$ contains a pseudo-loop modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$.
-In all the other cases, we can directly apply Lemma 3.5, yielding a pseudo-loop modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ within $E$.
+In the first case, let $S\subseteq E$ be smooth, weakly connected, and of pseudo-algebraic length $1$ modulo $\operatorname{Aut}(\mathbb{Q})$.
+By Lemma 3.9, $\langle{S^{\prime}}\rangle_{\ell\ell}\subseteq E$ contains a pseudo-loop modulo $\operatorname{Aut}(\mathbb{Q})$.
+In all the other cases, we can directly apply Lemma 3.5, yielding a pseudo-loop modulo $\operatorname{Aut}(\mathbb{Q})$ within $E$.
 ∎
-Similarly, putting together Lemmata 3.10 and 3.5, we obtain the corresponding pseudo-loop lemma for symmetric hypergraphs $\mathbb{T}$.
-By what we have shown, it in fact suffices for $\mathbb{T}$ to be cyclic and $2$-transitive.
-This amounts to a strictly weaker condition than full symmetry: for example, for every odd number $n\geq 7$, the alternating group $A_{n}$ is a proper subgroup of $\textnormal{Sym}(n)$ with these properties.
+Similarly, putting together Lemmata 3.11 and 3.5, we obtain the corresponding pseudo-loop lemma for cyclic and $2$-transitive hypergraphs $\mathbb{T}$.
 It might be of interest to point out that pseudo-loop lemmata involving symmetry conditions provided by non-trivial group actions have also been established for finite hypergraphs (see [BK12, Bru23]).
-See 1.6
+See 1.5
 5. Pseudo-loop conditions from pseudo-loops
-Following the standard method – established first by Siggers [Sig10] for finite structures, and adapted to $\omega$-categorical ones by Barto-Pinsker [BP20] – the existence of pseudo-loops in Theorems 1.3 and 1.6 implies the validity of the corresponding pseudo-loop conditions in all temporal constraint languages that are not omni-expressive.
+Following the standard method – established first by Siggers [Sig10] for finite structures, and adapted to $\omega$-categorical ones by Barto-Pinsker [BP20] – the existence of pseudo-loops in Theorems 1.3 and 1.5 implies the validity of the corresponding pseudo-loop conditions in all temporal constraint languages that are not omni-expressive.
+For the reader’s convenience, we provide the proof of Corollary 1.4.
 See 1.4
@@
 Suppose that $\mathbb{G}=(\{1,\dots,n\};E_{\mathbb{G}})$ is a finite smooth digraph of algebraic length $1$, and $u\circ s(x_{i_{1}},\dots,x_{i_{m}})\approx v\circ s(x_{j_{1}},\dots,x_{j_{m}})$ is the corresponding $\mathbb{G}$-pseudo-loop condition.
-We show that for every $k\geq 1$ and for every $a_{1},\dots,a_{n}\in\mathbb{Q}^{k}$ there exists an $m$-ary operation $s\in\operatorname{\textnormal{Pol}}(\mathbb{A})$ such that $s(a_{i_{1}},\dots,a_{i_{m}})$ and $s(a_{j_{1}},\dots,a_{j_{m}})$ are contained in the same $\operatorname{\textnormal{Aut}}(\mathbb{Q})$-orbit.
-A standard compactness argument (as provided, for example, in [BP20, Lemma 4.2]) using oligomorphicity of $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ then yields the ‘global’ validity in $\operatorname{\textnormal{Pol}}(\mathbb{A})$ of the desired pseudo-identity.
+We show that for every $k\geq 1$ and for every $a_{1},\dots,a_{n}\in\mathbb{Q}^{k}$ there exists an $m$-ary operation $s\in\operatorname{Pol}(\mathbb{A})$ such that $s(a_{i_{1}},\dots,a_{i_{m}})$ and $s(a_{j_{1}},\dots,a_{j_{m}})$ are contained in the same $\operatorname{Aut}(\mathbb{Q})$-orbit.
+A standard compactness argument (as provided, for example, in [BP20, Lemma 4.2]) using oligomorphicity of $\operatorname{Aut}(\mathbb{Q})$ then yields the ‘global’ validity in $\operatorname{Pol}(\mathbb{A})$ of the desired pseudo-identity.
 To this end, take $k\geq 1$, and let $a_{1},\dots,a_{n}$ be arbitrary $k$-tuples of elements of $\mathbb{Q}$.
-Consider the binary relation $R\subseteq\mathbb{Q}^{k}\times\mathbb{Q}^{k}$ that consists precisely of all the tuples $(a_{i},a_{j})$ for which $(i,j)\in E_{\mathbb{G}}$, and let $E_{\mathbb{H}}\coloneqq\langle R\rangle_{\operatorname{\textnormal{Pol}}(\mathbb{A})}$.
-Since $E_{\mathbb{H}}$ is, by construction, preserved by all polymorphisms of $\mathbb{A}$, it is pp-definable in $\mathbb{A}$ [BN06].
-It follows that $\mathbb{H}\coloneqq(A^{k};E_{\mathbb{H}})$ is pp-interpretable in $\mathbb{A}$.
+Consider the binary relation $R\subseteq\mathbb{Q}^{k}\times\mathbb{Q}^{k}$ that consists precisely of all the tuples $(a_{i},a_{j})$ for which $(i,j)\in E_{\mathbb{G}}$, and let $E_{\mathbb{H}}$ be the smallest binary relation on $\mathbb{Q}^{k}$ that contains $R$ and is preserved by $\operatorname{Pol}(\mathbb{A})$.
+Since $E_{\mathbb{H}}$ is preserved by all polymorphisms of $\mathbb{A}$, it is pp-definable in $\mathbb{A}$ [BN06].
+It follows that $\mathbb{H}\coloneqq(\mathbb{Q}^{k};E_{\mathbb{H}})$ is pp-interpretable in $\mathbb{A}$.
 Moreover, observe that $\mathbb{H}$ is a smooth digraph of algebraic length $1$.
-Applying Theorem 1.3, we get a pseudo-loop modulo $\operatorname{\textnormal{Aut}}(\mathbb{Q})$ in $\mathbb{H}$.
-Unravelling definitions, this means that there exists $s\in\operatorname{\textnormal{Pol}}(\mathbb{A})$ such that $s(a_{i_{1}},\dots,a_{i_{m}})$ and $s(a_{j_{1}},\dots,a_{j_{m}})$ are contained in the same $\operatorname{\textnormal{Aut}}(\mathbb{Q})$-orbit.
+Applying Theorem 1.3, we get a pseudo-loop modulo $\operatorname{Aut}(\mathbb{Q})$ in $\mathbb{H}$.
+Unravelling definitions, this means that there exists $s\in\operatorname{Pol}(\mathbb{A})$ such that $s(a_{i_{1}},\dots,a_{i_{m}})$ and $s(a_{j_{1}},\dots,a_{j_{m}})$ are contained in the same $\operatorname{Aut}(\mathbb{Q})$-orbit.
 ∎
-The corresponding statement for loop conditions induced by symmetric hypergraphs is derived from Theorem 1.6 in a similar way.
-Again, we may replace the word ‘symmetric’ by ‘cyclic and $2$-transitive’.
-See 1.7
 6. Min-clean tuples
-Recall the definition of a min-clean tuple: for an $n$-ary relation $R$ on $\mathbb{Q}^{k}$, we say that $t\in R$ is min-clean if $\operatorname{\textnormal{minx}}(t_{i})=\operatorname{\textnormal{minx}}(t_{j})$ for all $i,j\in M(t)$, where
+Recall the definition of a min-clean tuple: for an $n$-ary relation $R$ on $\mathbb{Q}^{k}$, we say that $t\in R$ is min-clean if $\operatorname{minx}(t_{i})=\operatorname{minx}(t_{j})$ for all $i,j\in M(t)$, where
 $M(t)\coloneqq\left\{i\in[n]\colon\min(t_{i})=\min(t)\right\}.$
@@
 Clearly, every tuple $t\in R$ that satisfies $|M(t)|=1$ is both min-clean and min-ready in $R$.
+Min-ready tuples turn out useful for proving the existence of min-clean tuples.
+6.1. Min-clean tuples for $\operatorname{mi}$
+The following technical lemma holds true for both digraphs and hypergraphs.
+Lemma 6.1.
+Let $n\geq 2$, and let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{mi}\rangle$.
+If $t^{1}\in R$ is min-ready, and $t^{2},\dots,t^{m}\in R$ satisfy $i\in M(t^{1})\cap\dots\cap M(t^{m}),$ $\operatorname{minx}(t^{1}_{i})\cap\dots\cap\operatorname{minx}(t^{m}_{i})\neq\varnothing$, $\min(t_{i}^{1})=\dots=\min(t_{i}^{m})$, and $j\in M(t^{1})$, then also $j\in M(t^{1})\cap\dots\cap M(t^{m})$ and $\operatorname{minx}(t^{1}_{j})\cap\dots\cap\operatorname{minx}(t^{m}_{j})\neq\varnothing.$
+Proof.
+Let $c=\min(t_{i}^{1})$, and choose mappings $\alpha,\beta$ and $\gamma$ satisfying the requirements in Definition 2.1 in a way such that $\alpha(c)=c$.
+Consider the tuple $u\coloneqq{\operatorname{mi}^{[m]}}(t^{1},\dots,t^{m})$.
+Note that $\operatorname{minx}(u_{i})=\operatorname{minx}(t_{i}^{1})\cap\dots\cap\operatorname{minx}(t_{i}^{m})$, and that $\min(u_{i})=c$.
+By min-readiness of $t^{1}$, also $\min(u_{j})=c$.
+As $\beta(c),\gamma(c)>\alpha(c)=c$, we must therefore have $j\in M(t^{1})\cap\dots\cap M(t^{m})$ and $\operatorname{minx}(t_{j}^{1})\cap\dots\cap\operatorname{minx}(t_{j}^{m})\neq\varnothing$.
+∎
+We now present the core construction for smooth digraphs of pseudo-algebraic length $1$.
+Lemma 6.2.
+Let $S$ be a binary relation on $\mathbb{Q}^{k}$.
+If $S$ is smooth, weakly connected, and of pseudo-algebraic length $1$, then either $\langle{S}\rangle_{\operatorname{mi}}$ contains a tuple $t$ with $|M(t)|=1$, or $\bigcap_{a\in\operatorname{supp}{S}}\operatorname{minx}(a)\neq\varnothing$.
+Proof.
+Let $R\coloneqq\langle{S}\rangle_{\operatorname{mi}}$, and suppose that $|M(t)|=2$ for all $t\in R$.
+In particular, all elements of $R$ are min-ready.
+Moreover, observe that $R$ is weakly connected.
+It follows that $\min(t_{1})=\min(t_{2})$ for all $t\in R$.
+We may therefore apply Lemma 6.1 to any tuples $t^{1},\dots,t^{m}\in R$ satisfying either $\operatorname{minx}(t^{1}_{1})\cap\dots\cap\operatorname{minx}(t^{m}_{1})\neq\varnothing$ or $\operatorname{minx}(t^{1}_{2})\cap\dots\cap\operatorname{minx}(t^{m}_{2})\neq\varnothing$.
+Now, take $a,b\in\operatorname{supp}(S)\subseteq\operatorname{supp}(R)$ arbitrarily.
+Since $R$ is smooth, weakly connected, and has pseudo-algebraic length $1$, its $\sim_{k}$-factor is $m$-linked for some $m\in\mathbb{N}$.
+Therefore, there exist $b^{\prime}\sim_{k}b$ and an $m$-fence in $R$ from $a$ to $b^{\prime}$, i.e.
+an $R$-walk of the form
+$(x_{1,0}Rx_{1,1}R\dots Rx_{1,m}=x_{2,m}R^{-1}x_{2,m-1}R^{-1}\dots R^{-1}x_{2,0}=x_{3,0}R\dots R^{-1}x_{n,1}R^{-1}x_{n,0}),$
+where $a=x_{1,0}$ and $b^{\prime}=x_{n,0}$.
+Since $b^{\prime}\sim_{k}b$, we have $\operatorname{minx}(b)=\operatorname{minx}(b^{\prime})$.
+We claim that there is a coordinate on which every lower tip of the above $m$-fence admits its minimal entry, that is
+$\operatorname{minx}(a)\cap\bigcap_{1<\ell<n}\operatorname{minx}(x_{\ell,0})\cap\operatorname{minx}(b^{\prime})\neq\varnothing.$ (7)
+Indeed, applying Lemma 6.1 to the tuples $\binom{x_{1,m-1}}{x_{1,m}}$ and $\binom{x_{2,m-1}}{x_{1,m}}$, we obtain $\operatorname{minx}(x_{1,m-1})\cap\operatorname{minx}(x_{2,m-1})\neq\varnothing$.
+Consequently, we also get $\operatorname{minx}(x_{1,m-2})\cap\operatorname{minx}(x_{2,m-2})\neq\varnothing$, and, after $m$ steps, $\operatorname{minx}(a)\cap\operatorname{minx}(x_{2,0})\neq\varnothing$.
+We now apply Lemma 6.1 to the tuples $\binom{a}{x_{1,1}},\binom{x_{2,0}}{x_{2,1}}$ and $\binom{x_{2,0}}{x_{3,1}}$ of $R$, and deduce that also $\operatorname{minx}(x_{1,m})\cap\operatorname{minx}(x_{3,m})\neq\varnothing$.
+Continuing like this, we ultimately see that Equation 7 holds true.
+Since the $\sim_{k}$-factor of $S$ is finite, we can select finitely many elements $a_{1},a_{2},\dots,a_{\ell}\in\operatorname{supp}(E)$ such that each $\sim_{k}$-class appearing in $\operatorname{supp}(S)$ is represented by some $a_{i}$.
+Moreover, we may choose these elements so that for every $i<\ell$, there exists an $m$-fence in $R$ connecting $a_{i}$ to $a_{i+1}$.
+By concatenating these fences, we obtain an $m$-fence whose set of lower tips contains all the elements $a_{1},a_{2},\dots,a_{\ell}$.
+Repeating the argument in the previous paragraph, we see that $\bigcap_{i\leq\ell}\operatorname{minx}(a_{i})\neq\varnothing$.
+By the choice of the $a_{i}$, this intersection is equal to $\bigcap_{a\in\operatorname{supp}(S)}\operatorname{minx}(a)$.
+∎
+The existence of min-clean tuples is an almost immediate consequence of Lemma 6.2.
+Lemma 6.3.
+Let $E$ be a binary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{mi}\rangle$. If $E$ is smooth and has pseudo-algebraic length 1, then $E$ contains a min-clean tuple.
+Proof.
+Any tuple $t\in E$ with $|M(t)|=1$ is min-clean.
+Suppose that such a tuple does not exist in $E$.
+Since $E$ is smooth and has pseudo-algebraic length $1$, we can pick $S\subseteq E$ that is smooth, weakly connected, and has pseudo-algebraic length $1$.
+It follows that all tuples in $\operatorname{supp}(S)$ share the same minimal entry $c$.
+Pick $t^{1},\dots,t^{m}$ such that every $\sim_{2k}$-class of a tuple from $S$ is represented by one of $t^{1},\dots,t^{m}$.
+By smoothness of $S$, every $\sim_{k}$-class appearing among $t^{1}_{1},\dots,t^{m}_{1}$ coincides with some $\sim_{k}$-class appearing among $t^{1}_{2},\dots,t^{m}_{2}$, and vice versa.
+In the definition of $\operatorname{mi}$, choose $\alpha,\beta$ and $\gamma$ such that $\alpha(c)=c$.
+Applying Lemma 6.2 to $S$, we see that $\bigcap_{i\in[m]}\operatorname{minx}(t^{i}_{j})=\bigcap_{a\in\operatorname{supp}(S)}\operatorname{minx}(x)\neq\varnothing$ for $j=1,2.$ The tuple $t\coloneqq{\operatorname{mi}^{[m]}}(t^{1},\dots,t^{m})$ now satisfies $\min(t_{1})=\min(t_{2})=c$, and $\operatorname{minx}(t_{1})=\bigcap_{a\in\operatorname{supp}(S)}\operatorname{minx}(a)=\operatorname{minx}(t_{2})$.
+∎
+The following lemma, in the setting of $2$-transitive higher-arity relations $R$, either directly yields a min-clean tuple, or implies that $R$ consists entirely of min-ready tuples.
+Lemma 6.4.
+Let $n\geq 3$, and let $R$ be an $n$-ary 22-transitive relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{mi}\rangle$.
+Either there is $t\in R$ with $|M(t)|=1,$ or $|M(t)|=n$ for all $t\in R$.
+Proof.
+Suppose that $t\in R$ satisfies $|M(t)|<n,$ and without loss of generality assume that $1\in M(t)$.
+By $2$-transitivity of $R$, there are tuples $t^{2},\dots,t^{n}\in R$ such that $1\in M(t^{i})$, $i\notin M(t^{i})$, and $\min(t^{i})=\min(t)$ for all $2\leq i\leq n$.
+In Definition 2.1, choose $\alpha$ in the definition of $\operatorname{mi}$ such that $\alpha(\min(t))=\min(t).$ It then follows that the tuple $s\coloneqq{\operatorname{mi}^{[n-1]}}(t^{2},\dots,t^{n})$ satisfies $\min(s)=\min(t)$, and $M(s)=\{1\}$.
+∎
+Finally, we derive min-clean tuples for $2$-transitive relations. We point out that the reasoning in the first paragraph of the proof crucially relies on the assumption $n\geq 3$.
+Lemma 6.5.
+Let $n\geq 3$, and let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{mi}\rangle$. If $R$ is cyclic and $2$-transitive, then $R$ contains a min-clean tuple.
+Proof.
+Suppose $\min_{t\in R}|M(t)|>1$, and take $t\in R$ arbitrarily.
+By Lemma 6.4, $t$ is in particular min-ready, and it satisfies $\min(t_{1})=\dots=\min(t_{n})=:c$.
+We claim that $\bigcap_{i\leq n}\operatorname{minx}(t_{i})\neq\varnothing.$ Indeed, define $t^{1}\coloneqq t$, and take $t^{2},\dots,t^{n-1}\in R$ to be permutations of $t$ satisfying $t^{i}_{1}=t_{1}$ and $t^{i}_{2}=i+1$ for all $2\leq i\leq n-1$.
+This is possible by $2$-transitivity of $R$.
+Moreover, we may choose $t^{n}\in R$ that is a permutation of $t$ and satisfies $t^{n}_{2}=t_{1}$ and $t^{n}_{3}=t_{3}$.
+Since $\bigcap_{i\leq n-1}\operatorname{minx}(t_{1}^{i})=\operatorname{minx}(t_{1})$, Lemma 6.1 applied to the tuples $t^{1},\dots,t^{n-1}$ gives $\bigcap_{i\leq n-1}\operatorname{minx}(t^{i}_{j})\neq\varnothing$ for all $j\geq 2$.
+As, by definition, it holds that $\operatorname{minx}(t^{n}_{3})=\operatorname{minx}(t^{1}_{3})$ and hence $\bigcap_{i\leq n}\operatorname{minx}(t^{i}_{3})\neq\varnothing$, we may apply Lemma 6.1 again to the tuples $t^{1},\dots,t^{n}$.
+We obtain $\bigcap_{i\leq n}\operatorname{minx}(t^{i}_{j})\neq\varnothing$ for all $j\leq n$.
+By the choice of $t^{i}$, this intersection is for $j=2$ equal to $\bigcap_{i\geq 2}\operatorname{minx}(t_{i})$.
+Let now $s^{1},\dots,s^{n}$ denote all cyclic permutations of the tuple $t$.
+Since $R$ is cyclic, we have $s^{i}\in R$ for every $i\leq n$.
+Choose $\alpha$ as in Definition 2.1 such that $\alpha(c)=c,$ and set $s\coloneqq{\operatorname{mi}^{[n]}}(s^{1},\dots,s^{n}).$ Observe that $\min(s)=c$, and $\operatorname{minx}(s_{i})=\bigcap_{j\leq n}\operatorname{minx}(t_{j})$ for all $i\leq n$.
+In particular, $s$ is min-clean.
+∎
 6.2. Min-clean tuples for $\min$
 Lemma 6.6.
-Let $E$ be a binary relation on $\mathbb{Q}^{k}$. If $E$ is smooth and preserved by $\min$, then $E$ contains a min-clean tuple.
+Let $E$ be a binary relation on $\mathbb{Q}^{k}$ that is preserved by $\min$. If $E$ is smooth, then $E$ contains a min-clean tuple.
 Proof.
 If $E$ contains a tuple $t$ with $|M(t)|=1,$ then $t$ is min-clean.
-Otherwise, take $W$ to be any weakly connected component of $E$.
-Observe that by connectivity of $E|_{W\times W}$, all vertices $a,b$ contained in $W$ share the same minimal entry.
-It follows that $\operatorname{\textnormal{minx}}(\min(a,b))=\operatorname{\textnormal{minx}}(a)\cup\operatorname{\textnormal{minx}}(b)$ for all $a,b\in W$.
-Pick $t^{1},\dots,t^{m}\in E|_{W\times W}$ whose $\sim_{2k}$-classes represent all $\sim_{2k}$-classes appearing in $E|_{W\times W}$.
+Otherwise, take $S\subseteq E$ to be smooth and weakly connected.
+Observe that by connectivity of $S$, all vertices $a,b$ contained in $\operatorname{supp}(S)$ share the same minimal entry.
+It follows that $\operatorname{minx}(\min(a,b))=\operatorname{minx}(a)\cup\operatorname{minx}(b)$ for all $a,b\in\operatorname{supp}(S)$.
+Pick $t^{1},\dots,t^{m}\in S$ whose $\sim_{2k}$-classes represent all $\sim_{2k}$-classes appearing in $S$.
 Clearly, this is possible as there are only finitely many $\sim_{2k}$-classes on $\mathbb{Q}^{2k}$.
-Observe that by smoothness of $E|_{W\times W}$, the set of $\sim_{k}$-classes appearing among $t_{1}^{1},\dots,t_{1}^{m}$ coincides with the set of $\sim_{k}$-classes of $t_{2}^{1},\dots,t_{2}^{m}$.
-Therefore, the tuple $t\coloneqq{\min^{[m]}}(t^{1},\dots,t^{m})$ is min-clean because it satisfies $\operatorname{\textnormal{minx}}(t_{1})=\bigcup_{a\in W}\operatorname{\textnormal{minx}}(a)=\operatorname{\textnormal{minx}}(t_{2})$.
+Observe that by smoothness of $S$, the set of $\sim_{k}$-classes appearing among $t_{1}^{1},\dots,t_{1}^{m}$ coincides with the set of $\sim_{k}$-classes of $t_{2}^{1},\dots,t_{2}^{m}$.
+Therefore, the tuple $t\coloneqq{\min^{[m]}}(t^{1},\dots,t^{m})$ is min-clean because it satisfies $\operatorname{minx}(t_{1})=\bigcup_{a\in\operatorname{supp}(S)}\operatorname{minx}(a)=\operatorname{minx}(t_{2})$.
 ∎
-As it turns out, in the case of symmetric relations preserved by $\min$, our endeavour to prove the existence of min-clean tuples leads us to the immediate conclusion that every such relation must contain an actual loop.
-In fact, it even suffices that the relation is cyclic.
+As it turns out, in the case of cyclic relations preserved by $\min$, our endeavour to prove the existence of min-clean tuples leads us to the immediate conclusion that every such relation must contain an actual loop.
 Lemma 6.7.
-Let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$. If $R$ is cyclic and preserved by $\min$, then $R$ contains a loop.
+Let $n\geq 2$, and let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\min$. If $R$ is cyclic, then $R$ contains a loop.
 Proof.
@@
 ∎
-5.2. Min-clean tuples for $\operatorname{mi}$
+6.3. Min-clean tuples for $\operatorname{mx}$
 Lemma 6.8.
-Let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\operatorname{mi}$, $t^{1}\in R$ min-ready, and $t^{2},\dots,t^{m}\in R$.
-If $i\in M(t^{1})\cap\dots\cap M(t^{m}),$ $\operatorname{\textnormal{minx}}(t^{1}_{i})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{i})\neq\varnothing$, $\min(t_{i}^{1})=\dots=\min(t_{i}^{m})$, and $j\in M(t^{1})$, then also $j\in M(t^{1})\cap\dots\cap M(t^{m})$ and $\operatorname{\textnormal{minx}}(t^{1}_{j})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{j})\neq\varnothing.$
+Let $E$ be a binary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{mx}\rangle$.
+Either $E$ contains a tuple $t$ with $|M(t)|=1$, or for all $t,t^{\prime}\in E$ with $\min(t_{i})=\min(t^{\prime}_{i})$ and $\operatorname{minx}(t_{i})=\operatorname{minx}(t^{\prime}_{i})$ for $i\in\{1,2\}$, also $\operatorname{minx}(t_{j})=\operatorname{minx}(t^{\prime}_{j})$ for $j\neq i$.
 Proof.
-Let $c=\min(t_{i}^{1})$, and choose mappings $\alpha,\beta$ and $\gamma$ satisfying the requirements in 2.1 in a way such that $\alpha(c)=c$.
-Consider the tuple $u\coloneqq{\operatorname{mi}^{[m]}}(t^{1},\dots,t^{m})$.
-Note that $\operatorname{\textnormal{minx}}(u_{i})=\operatorname{\textnormal{minx}}(t_{i}^{1})\cap\dots\cap\operatorname{\textnormal{minx}}(t_{i}^{m})$, and that $\min(u_{i})=c$.
-By min-readiness of $t^{1}$, also $\min(u_{j})=c$.
-As $\beta(c),\gamma(c)>\alpha(c)=c$, we must therefore have $j\in M(t^{1})\cap\dots\cap M(t^{m})$ and $\operatorname{\textnormal{minx}}(t_{j}^{1})\cap\dots\cap\operatorname{\textnormal{minx}}(t_{j}^{m})\neq\varnothing$.
+Suppose that all tuples $t$ satisfy $|M(t)|=2$.
+Without loss of generality, let $t,t^{\prime}\in E$ be such that $c\coloneqq\min(t_{1})=\min(t^{\prime}_{1})$ and $\operatorname{minx}(t_{1})=\operatorname{minx}(t^{\prime}_{1})$.
+Observe that for all $a,b\in\mathbb{Q}^{k}$ with $\min(a)=\min(b)$ it holds that $\min(\operatorname{mx}(a,b))=\beta(\min(a))$ if and only if $\operatorname{minx}(a)\triangle\operatorname{minx}(b)=\varnothing$, where $\beta$ is the endomorphism in the definition of $\operatorname{mx}$.
+The tuple $s\coloneqq\operatorname{mx}(t,t^{\prime})$ now satisfies $\min(s_{1})=\beta(c)$.
+By assumption, we have $\min(s_{2})=\min(s_{1})=\beta(c)$, and thus indeed $\operatorname{minx}(t_{2})\triangle\operatorname{minx}(t_{2}^{\prime})=\varnothing$.
 ∎
 Lemma 6.9.
-Let $E$ be a binary relation on $\mathbb{Q}^{k}$.
-Suppose that $W$ is a weakly connected component of $E$ of pseudo-algebraic length $1$.
-If $E$ is smooth and preserved by $\operatorname{mi}$, then either $E$ contains a tuple $t$ with $|M(t)|=1$, or $\bigcap_{a\in W}\operatorname{\textnormal{minx}}(a)\neq\varnothing$.
-Proof.
-If $|M(t)|=2$ for all $t\in E$, then by connectivity of $E|_{W\times W}$ we have $\min(a)=\min(b)$ for all vertices $a,b$ appearing in $W$.
-Take $a,b\in W$ arbitrarily.
-Since $E$ has pseudo-algebraic length $1$, the $\sim_{k}$-factor of the relation $E|_{W\times W}$ is smooth and $m$-linked for some $m\in\mathbb{N}$.
-Therefore, there exist $b^{\prime}\sim_{k}b$ and an $m$-fence in $E$ from $a$ to $b^{\prime}$, i.e.
-an $E$-walk of the form
-$(x_{1,0}Ex_{1,1}E\dots Ex_{1,m}=x_{2,m}E^{-1}x_{2,m-1}E^{-1}\dots E^{-1}x_{2,0}=x_{3,0}E\dots E^{-1}x_{n,1}E^{-1}x_{n,0}),$
-where $a=x_{1,0}$ and $b^{\prime}=x_{n,0}$.
-Since $b^{\prime}\sim_{k}b$, we have $\operatorname{\textnormal{minx}}(b)=\operatorname{\textnormal{minx}}(b^{\prime})$.
-We claim that there is a coordinate on which every lower tip of the above $m$-fence admits its minimal entry, that is
-$\operatorname{\textnormal{minx}}(a)\cap\bigcap_{\ell<n}\operatorname{\textnormal{minx}}(x_{\ell,0})\cap\operatorname{\textnormal{minx}}(b^{\prime})\neq\varnothing.$ (7)
-Indeed, applying Lemma 5.3 to the tuples $\binom{x_{1,m-1}}{x_{1,m}}$ and $\binom{x_{2,m-1}}{x_{1,m}}$, we obtain $\operatorname{\textnormal{minx}}(x_{1,m-1})\cap\operatorname{\textnormal{minx}}(x_{2,m-1})\neq\varnothing$.
-Consequently, we also get $\operatorname{\textnormal{minx}}(x_{1,m-2})\cap\operatorname{\textnormal{minx}}(x_{2,m-2})\neq\varnothing$, and, after $m$ steps, $\operatorname{\textnormal{minx}}(a)\cap\operatorname{\textnormal{minx}}(x_{2,0})\neq\varnothing$.
-We now apply Lemma 5.3 to the edges $\binom{a}{x_{1,1}},\binom{x_{2,0}}{x_{2,1}}$ and $\binom{x_{2,0}}{x_{3,1}}$, and deduce that also $\operatorname{\textnormal{minx}}(x_{1,m})\cap\operatorname{\textnormal{minx}}(x_{3,m})\neq\varnothing$.
-Continuing like this, we ultimately see that Equation 7 holds true.
-Since the $\sim_{k}$-factor of $W$ is finite, we can select finitely many elements $a_{1},a_{2},\dots,a_{\ell}\in W$ such that each $\sim_{k}$-class appearing in $W$ is represented by some $a_{i}$.
-Moreover, we may choose these elements so that for every $i<\ell$, there exists an $m$-fence in $E$ connecting $a_{i}$ to $a_{i+1}$.
-By concatenating these fences, we obtain an $m$-fence whose set of lower tips contains all the elements $a_{1},a_{2},\dots,a_{\ell}$.
-Repeating the argument in the previous paragraph, we see that $\bigcap_{i\leq\ell}\operatorname{\textnormal{minx}}(a_{i})\neq\varnothing$.
-By the choice of the $a_{i}$, this intersection is equal to $\bigcap_{a\in W}\operatorname{\textnormal{minx}}(a)$.
-∎
-Lemma 5.5.
-Let $E$ be a binary relation on $\mathbb{Q}^{k}$. If $E$ is smooth, has pseudo-algebraic length 1, and is preserved by $\operatorname{mi}$, then $E$ contains a min-clean tuple.
-Proof.
-Any tuple $t\in E$ with $|M(t)|=1$ is min-clean.
-Suppose that such a tuple does not exist in $E$, and fix a weakly connected component $W$ of $E$ that has pseudo-algebraic length $1$.
-It follows that all tuples in $W$ share the same minimal entry $c$.
-Pick $t^{1},\dots,t^{m}$ such that every $\sim_{2k}$-class of a tuple from $E|_{W\times W}$ is represented by one of $t^{1},\dots,t^{m}$.
-By smoothness of $E|_{W\times W}$, every $\sim_{k}$-class appearing among $t^{1}_{1},\dots,t^{m}_{1}$ coincides with some $\sim_{k}$-class appearing among $t^{1}_{2},\dots,t^{m}_{2}$, and vice versa.
-In the definition of $\operatorname{mi}$, choose $\alpha,\beta$ and $\gamma$ such that $\alpha(c)=c$.
-By Lemma 5.4, we have $\bigcap_{i\in[m]}\operatorname{\textnormal{minx}}(t^{i}_{j})=\bigcap_{a\in W}\operatorname{\textnormal{minx}}(x)\neq\varnothing$ for $j=1,2.$ The tuple $t\coloneqq{\operatorname{mi}^{[m]}}(t^{1},\dots,t^{m})$ now satisfies $\min(t_{1})=\min(t_{2})=c$, and $\operatorname{\textnormal{minx}}(t_{1})=\bigcap_{a\in W}\operatorname{\textnormal{minx}}(a)=\operatorname{\textnormal{minx}}(t_{2})$.
-∎
-Lemma 5.6.
-Let $R$ be an $n$-ary 22-transitive relation on $\mathbb{Q}^{k}$ that is preserved by $\operatorname{mi}$. Either there is $t\in R$ with $|M(t)|=1,$ or $|M(t)|=n$ for all $t\in R$.
-Proof.
-Suppose that $t\in R$ satisfies $|M(t)|<n,$ and without loss of generality assume that $1\in M(t)$.
-By $2$-transitivity of $R$, there are tuples $t^{2},\dots,t^{n}\in R$ such that $1\in M(t^{i})$, $i\notin M(t^{i})$, and $\min(t^{i})=\min(t)$ for all $2\leq i\leq n$.
-In 2.1, choose $\alpha$ in the definition of $\operatorname{mi}$ such that $\alpha(\min(t))=\min(t).$ It then follows that the tuple $s\coloneqq{\operatorname{mi}^{[n-1]}}(t^{2},\dots,t^{n})$ satisfies $\min(s)=\min(t)$, and $M(s)=\{1\}$.
-∎
-Lemma 5.7.
-Let $R$ be an $n$-ary cyclic and $2$-transitive relation on $\mathbb{Q}^{k}$ with $n\geq 3$. If $R$ is preserved by $\operatorname{mi}$, then $R$ contains a min-clean tuple.
-Proof.
-Suppose $\min_{t\in R}|M(t)|>1$, and take $t\in R$ arbitrarily.
-By Lemma 5.6, $t$ is min-ready and satisfies $\min(t_{1})=\dots=\min(t_{n})=:c$.
-We claim that $\bigcap_{i\leq n}\operatorname{\textnormal{minx}}(t_{i})\neq\varnothing.$ Indeed, define $t^{1}\coloneqq t$, and take $t^{2},\dots,t^{n-1}\in R$ to be permutations of $t$ satisfying $t^{i}_{1}=t_{1}$ and $t^{i}_{2}=i+1$ for all $2\leq i\leq n-1$.
-This is possible by $2$-transitivity of $R$.
-Moreover, we may choose $t^{n}\in R$ that is a permutation of $t$ and satisfies $t^{n}_{2}=t_{1}$ and $t^{n}_{3}=t_{3}$.
-Since $\bigcap_{i\leq n-1}\operatorname{\textnormal{minx}}(t_{1}^{i})=\operatorname{\textnormal{minx}}(t_{1})$, Lemma 5.3 applied to the tuples $t^{1},\dots,t^{n-1}$ gives $\bigcap_{i\leq n-1}\operatorname{\textnormal{minx}}(t^{i}_{j})\neq\varnothing$ for all $j\geq 2$.
-As, by definition, it holds that $\operatorname{\textnormal{minx}}(t^{n}_{3})=\operatorname{\textnormal{minx}}(t^{1}_{3})$ and hence $\bigcap_{i\leq n}\operatorname{\textnormal{minx}}(t^{i}_{3})\neq\varnothing$, we may apply Lemma 5.3 again to the tuples $t^{1},\dots,t^{n}$.
-We obtain $\bigcap_{i\leq n}\operatorname{\textnormal{minx}}(t^{i}_{j})\neq\varnothing$ for all $j\leq n$.
-By the choice of $t^{i}$, this intersection is for $j=2$ equal to $\bigcap_{i\geq 2}\operatorname{\textnormal{minx}}(t_{i})$.
-Let now $s^{1},\dots,s^{n}$ denote all cyclic permutations of the tuple $t$.
-Since $R$ is cyclic, we have $s^{i}\in R$ for every $i\leq n$.
-Choose $\alpha$ as in 2.1 such that $\alpha(c)=c,$ and set $s\coloneqq{\operatorname{mi}^{[n]}}(s^{1},\dots,s^{n}).$ Observe that $\min(s)=c$, and $\operatorname{\textnormal{minx}}(s_{i})=\bigcap_{j\leq n}\operatorname{\textnormal{minx}}(t_{j})$ for all $i\leq n$.
-In particular, $s$ is min-clean.
-∎
-5.3. Min-clean tuples for $\operatorname{mx}$
-Lemma 5.8.
-Let $E$ be a binary relation on $\mathbb{Q}^{k}$.
-If $E$ is preserved by $\operatorname{mx}$, then either $E$ contains a tuple $t$ with $|M(t)|=1$, or for all $t,t^{\prime}\in E$ with $\min(t_{i})=\min(t^{\prime}_{i})$ and $\operatorname{\textnormal{minx}}(t_{i})=\operatorname{\textnormal{minx}}(t^{\prime}_{i})$ for $i\in\{1,2\}$, also $\operatorname{\textnormal{minx}}(t_{j})=\operatorname{\textnormal{minx}}(t^{\prime}_{j})$ for $j\neq i$.
-Proof.
-Suppose that all tuples $t$ satisfy $|M(t)|=2$.
-Without loss of generality, let $t,t^{\prime}\in E$ be such that $c\coloneqq\min(t_{1})=\min(t^{\prime}_{1})$ and $\operatorname{\textnormal{minx}}(t_{1})=\operatorname{\textnormal{minx}}(t^{\prime}_{1})$.
-Observe that for all $a,b\in\mathbb{Q}^{k}$ with $\min(a)=\min(b)$ it holds that $\min(\operatorname{mx}(a,b))=\beta(\min(a))$ if and only if $\operatorname{\textnormal{minx}}(a)\triangle\operatorname{\textnormal{minx}}(b)=\varnothing$, where $\beta$ is the endomorphism in the definition of $\operatorname{mx}$.
-The tuple $s\coloneqq\operatorname{mx}(t,t^{\prime})$ now satisfies $\min(s_{1})=\beta(c)$.
-By assumption, we have $\min(s_{2})=\min(s_{1})=\beta(c)$, and thus indeed $\operatorname{\textnormal{minx}}(t_{2})\triangle\operatorname{\textnormal{minx}}(t_{2}^{\prime})=\varnothing$.
-∎
-Lemma 5.9.
-Let $E$ be a binary relation on $\mathbb{Q}^{k}$. If $E$ has pseudo-algebraic length $1$ and is preserved by $\operatorname{mx}$, then $E$ contains a min-clean tuple.
+Let $E$ be a binary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{mx}\rangle$. If $E$ is smooth and of pseudo-algebraic length $1$, then $E$ contains a min-clean tuple.
 Proof.
@@
 Suppose that $|M(t)|=2$ for all $t\in E$.
-Take $W$ a weakly connected component of pseudo-algebraic length $1$.
-As the $\sim_{k}$-factor of the relation $E|_{W\times W}$ is finite, smooth, weakly connected, and of algebraic length $1$, it is $m$-linked for some $m\in\mathbb{N}$.
-We claim that every $t\in E|_{W\times W}$ is min-clean.
-Indeed, for all $t\in E|_{W\times W}$ there are $t^{\prime}_{2}\sim_{k}t_{2}$ and an $m$-fence in $E|_{W\times W}$ connecting $t_{1}$ and $t^{\prime}_{2}$.
-Similar to the proof of Lemma 5.4, repeatedly applying Lemma 5.8 instead of Lemma 5.3, it follows that $\operatorname{\textnormal{minx}}(t_{1})=\operatorname{\textnormal{minx}}(t_{2})$.
+Take $S\subseteq E$ that is smooth, weakly connected, and has pseudo-algebraic length $1$.
+As the $\sim_{k}$-factor of $S$ is finite, smooth, weakly connected, and of algebraic length $1$, it is $m$-linked for some $m\in\mathbb{N}$.
+We claim that every $t\in S$ is min-clean.
+Indeed, for all $t\in S$ there are $t^{\prime}_{2}\sim_{k}t_{2}$ and an $m$-fence in $S$ connecting $t_{1}$ and $t^{\prime}_{2}$.
+Similar to the proof of Lemma 6.2, repeatedly applying Lemma 6.8 instead of Lemma 6.1 to the relation $\langle{S}\rangle_{\operatorname{mx}}\subseteq E$, it follows that $\operatorname{minx}(t_{1})=\operatorname{minx}(t_{2})$.
 Lemma 6.10.
-Let $R$ be an $n$-ary 22-transitive relation on $\mathbb{Q}^{k}$ with $n\geq 3$. If $R$ is preserved by $\operatorname{mx}$, then $R$ contains a min-clean tuple.
+Let $n\geq 3$, and let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{mx}\rangle$. If $R$ is cyclic, then $R$ contains a min-clean tuple.
 Proof.
-Suppose that $t\in R$ is not min-clean.
-In particular, we must have $|M(t)|>1$.
-Without loss of generality, assume $1,2\in M(t)$ and $\operatorname{\textnormal{minx}}(t_{1})\neq\operatorname{\textnormal{minx}}(t_{2})$.
-Observe that $\operatorname{\textnormal{minx}}(\operatorname{mx}(t_{1},t_{2}))=\operatorname{\textnormal{minx}}(t_{1})\triangle\operatorname{\textnormal{minx}}(t_{2})$, and $\min(\operatorname{mx}(t_{1},t_{2}))=\alpha(\min(t))$.
-By $2$-transitivity of $R$, there is a tuple $t^{\prime}\in R$ for which $t^{\prime}_{1}=t_{2}$ and $t^{\prime}_{2}=t_{1}$.
-The tuple $s\coloneqq\operatorname{mx}(t,t^{\prime})$ is contained in $R$, and it satisfies $M(s)=\{1,2\}$, $\min(s)=\alpha(\min t)$, and $\operatorname{\textnormal{minx}}(\textnormal{pr}_{1}(s))=\operatorname{\textnormal{minx}}(\textnormal{pr}_{2}(s))$.
+Let $t\in R$, and set $c\coloneqq\min(t)$.
+We choose $\alpha$ and $\beta$ in the definition of $\operatorname{mx}$ such that $\alpha$ fixes $c$, such that $\beta(c)$ is strictly smaller than the second smallest entry of $t$ (if such an entry exists), and such that $\alpha(\beta(c))=\beta(c)$.
+Denote the cyclic permutations of $t$ by $t^{1},\dots,t^{n}$, and set $s\coloneqq{\operatorname{mx}^{[n]}}(t^{1},\dots,t^{n})$.
+We claim that $s\in R$ is min-clean.
+To see this, for every $j\in[k]$ set
+$N(j)\coloneqq|\{i\in[n]:\operatorname{pr}_{j}(t_{i})=c\}|.$
+Observe that whenever $N(j)\neq 0$, then for all $i\in[n]$ the value of $\operatorname{pr}_{j}(s_{i})$ is either $c$ or $\beta(c)$, depending only on the parity of $N(j)$.
+Moreover, all remaining entries of $s$ are greater than $\beta(c)$.
+It follows that $s$ is indeed min-clean.
 ∎
 6.4. Min-clean tuples for $\ell\ell$
+Remarkably, for the existence of min-clean tuples in relations preserved by $\ell\ell$, we will in fact only make use of the binary operation $\operatorname{lex}$ that is generated by $\ell\ell$.
 Lemma 6.11.
-Let $R$ be preserved by $\operatorname{\textnormal{lex}}$, $t^{1}\in R$ min-ready, and $t^{2},\dots,t^{m}\in R$.
-If $i\in M(t^{1})\cap\dots\cap M(t^{m}),$ $\operatorname{\textnormal{minx}}(t^{1}_{i})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{i})\neq\varnothing$, and $j\in M(t_{1})$, then also $j\in M(t^{2})\cap\dots\cap M(t^{m})$ and $\operatorname{\textnormal{minx}}(t^{1}_{j})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{j})\neq\varnothing.$
+Let $n\geq 2$, and let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\operatorname{lex}$.
+If $t^{1}\in R$ is min-ready, and $t^{2},\dots,t^{m}\in R$ satisfy $i\in M(t^{1})\cap\dots\cap M(t^{m}),$ $\operatorname{minx}(t^{1}_{i})\cap\dots\cap\operatorname{minx}(t^{m}_{i})\neq\varnothing$, and $j\in M(t_{1})$, then also $j\in M(t^{2})\cap\dots\cap M(t^{m})$ and $\operatorname{minx}(t^{1}_{j})\cap\dots\cap\operatorname{minx}(t^{m}_{j})\neq\varnothing.$
 Proof.
-Consider the tuple $t$ defined by $t\coloneqq{\operatorname{\textnormal{lex}}^{[m]}}(t^{1},\dots,t^{m})$.
-Observe that $\operatorname{\textnormal{minx}}(t_{i})=\operatorname{\textnormal{minx}}(t^{1}_{i})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{i})$ and $\min(t)={\operatorname{\textnormal{lex}}^{[m]}}(\min(t^{1}_{i}),\dots,\min(t^{m}_{i}))$.
-By min-readiness of $t^{1}$, we also have $\min(t_{i})=\min(t_{j}),$ which by injectivity of $\operatorname{\textnormal{lex}}^{[m]}$} implies $\operatorname{\textnormal{minx}}(t^{1}_{j})\cap\dots\cap\operatorname{\textnormal{minx}}(t^{m}_{j})\neq\varnothing.$
+Consider the tuple $t$ defined by $t\coloneqq{\operatorname{lex}^{[m]}}(t^{1},\dots,t^{m})$.
+Observe that $\operatorname{minx}(t_{i})=\operatorname{minx}(t^{1}_{i})\cap\dots\cap\operatorname{minx}(t^{m}_{i})$ and $\min(t)={\operatorname{lex}^{[m]}}(\min(t^{1}_{i}),\dots,\min(t^{m}_{i}))$.
+By min-readiness of $t^{1}$, we also have $\min(t_{i})=\min(t_{j}),$ which by injectivity of $\operatorname{lex}^{[m]}$} implies $\operatorname{minx}(t^{1}_{j})\cap\dots\cap\operatorname{minx}(t^{m}_{j})\neq\varnothing.$
 Lemma 6.12.
-Let $E$ be a binary relation on $\mathbb{Q}^{k}$.
-Suppose that $W$ is a weakly connected component of $E$ of pseudo-algebraic length $1$.
-If $E$ is smooth and preserved by $\operatorname{\textnormal{lex}}$, then either $E$ contains a tuple $t$ with $|M(t)|=1$, or $\bigcap_{a\in W}\operatorname{\textnormal{minx}}(a)\neq\varnothing$.
+Let $S$ be a binary relation on $\mathbb{Q}^{k}$.
+If $S$ is smooth, weakly connected, and of pseudo-algebraic length $1$, then either $\langle{S}\rangle_{\operatorname{lex}}$ contains a tuple $t$ with $|M(t)|=1$, or $\bigcap_{a\in\operatorname{supp}(S)}\operatorname{minx}(a)\neq\varnothing$.
 Proof.
@@
 Lemma 6.13.
-Let $E$ be a binary relation on $\mathbb{Q}^{k}$ that is smooth and preserved by $\operatorname{\textnormal{lex}}$.
-If $S\subseteq E$ is smooth itself, has pseudo-algebraic length $1$, and $\textnormal{pr}_{1}(S)$ is weakly connected, then $S^{\prime}$ contains a min-clean tuple.
+Let $S$ be a binary relation on $\mathbb{Q}^{k}$. If $S$ is smooth, weakly connected, and of pseudo-algebraic length $1$, then $S^{\prime}$ contains a min-clean tuple.
 Proof.
 By Lemma 3.7, $S^{\prime}$ is non-empty.
-If $E$ contains a tuple $t$ with $|M(t)|=1$, then for every $s\in S^{\prime}$ the tuple $\operatorname{\textnormal{lex}}(t,s)$ is min-clean and contained in $S^{\prime}$.
-Suppose now that $|M(t)|=2$ for all $t\in E$.
+If $\langle{S}\rangle_{\operatorname{lex}}$ contains a tuple $t$ with $|M(t)|=1$, then for every $s\in S^{\prime}$ the tuple $\operatorname{lex}(t,s)$ is min-clean and contained in $S^{\prime}$.
+Suppose now that $|M(t)|=2$ for all $t\in\langle{S}\rangle_{\operatorname{lex}}$.
+In particular, every element of $\langle{S}\rangle_{\operatorname{lex}}$ is min-ready.
 By connectedness, all tuples in $S$ share the same minimal entry $c$.
@@
 Thus, by the choice of $t^{1},\dots,t^{m}$, we have
-$\bigcap_{i\in[m]}\operatorname{\textnormal{minx}}(t_{1}^{i})=\bigcap_{i\in[m]}\operatorname{\textnormal{minx}}(t_{2}^{i})=\bigcap_{a\in\textnormal{pr}_{1}(S)}\operatorname{\textnormal{minx}}(a).$
-Lemma 5.12 applied to the weakly connected component of $E$ containing $\textnormal{pr}_{1}(S)$ yields $\bigcap_{a\in\textnormal{pr}_{1}(S)}\operatorname{\textnormal{minx}}(a)\neq\varnothing.$ Therefore, the tuple $s\coloneqq{\operatorname{\textnormal{lex}}^{[m]}}(t^{1},\dots,t^{m})$ satisfies $\min(s_{1})=\min(s_{2})={\operatorname{\textnormal{lex}}^{[m]}}(c,\dots,c)$, as well as $\operatorname{\textnormal{minx}}(s_{1})=\bigcap_{a\in\textnormal{pr}_{1}(S)}\operatorname{\textnormal{minx}}(a)=\operatorname{\textnormal{minx}}(s_{2})$.
-Moreover, by the choice of $t^{1},\dots,t^{m}$ and injectivity of $\operatorname{\textnormal{lex}}^{[m]}$}, we also have $s\in S^{\prime}$.
+$\bigcap_{i\in[m]}\operatorname{minx}(t_{1}^{i})=\bigcap_{i\in[m]}\operatorname{minx}(t_{2}^{i})=\bigcap_{a\in\operatorname{supp}(S)}\operatorname{minx}(a).$
+Lemma 6.12 yields $\bigcap_{a\in\operatorname{supp}(S)}\operatorname{minx}(a)\neq\varnothing$.
+Therefore, the tuple $s\coloneqq{\operatorname{lex}^{[m]}}(t^{1},\dots,t^{m})$ satisfies $\min(s_{1})=\min(s_{2})={\operatorname{lex}^{[m]}}(c,\dots,c)$, as well as $\operatorname{minx}(s_{1})=\bigcap_{a\in\operatorname{supp}(S)}\operatorname{minx}(a)=\operatorname{minx}(s_{2})$.
+Moreover, by the choice of $t^{1},\dots,t^{m}$ and injectivity of $\operatorname{lex}^{[m]}$}, we also have $s\in S^{\prime}$.
 ∎
 Lemma 6.14.
-Let $R$ be an $n$-ary 22-transitive relation on $\mathbb{Q}^{k}$ that is preserved by $\operatorname{\textnormal{lex}}$. Either there is $t\in R$ with $|M(t)|=1,$ or $|M(t)|=n$ for all $t\in R$.
+Let $n\geq 2$, and let $R$ be an $n$-ary 22-transitive relation on $\mathbb{Q}^{k}$ that is preserved by $\operatorname{lex}$. Either there is $t\in R$ with $|M(t)|=1,$ or $|M(t)|=n$ for all $t\in R$.
 Proof.
-The proof is similar to the one of Lemma 5.6, using $\operatorname{\textnormal{lex}}^{[n-1]}$} in place of $\operatorname{mi}^{[n-1]}$}. ∎
+The proof is similar to the one of Lemma 6.4, using $\operatorname{lex}^{[n-1]}$} in place of $\operatorname{mi}^{[n-1]}$}. ∎
 Lemma 6.15.
-Let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ with $n\geq 3$. If $R$ is cyclic, $2$-transitive, and preserved by $\operatorname{\textnormal{lex}}$, then $R^{\prime}$ contains a min-clean tuple.
+Let $n\geq 3$, and let $R$ be an $n$-ary relation on $\mathbb{Q}^{k}$ that is preserved by $\langle\operatorname{lex}\rangle$.
+If $R$ is cyclic and $2$-transitive, then $R^{\prime}\subseteq R$ contains a min-clean tuple.
 Proof.
-If $R$ contains a tuple $t$ with $|M(t)|=1$, this tuple is min-clean.
-It then follows that for any $s\in R^{\prime}$ (which exists by Lemma 3.7), the tuple $\operatorname{\textnormal{lex}}(t,s)\in R^{\prime}$ is also min-clean.
-If $R$ does not contain such a tuple, then by Lemma 5.14 all tuples satisfy $|M(t)|=n$ and are, in particular, min-ready.
+If $R^{\prime}$ contains a tuple $t$ with $|M(t)|=1$, this tuple is min-clean.
+If $R^{\prime}$ (which is non-empty by Lemma 3.7) does not contain such a tuple, then by Lemma 6.14 all tuples in $R^{\prime}$ satisfy $|M(t)|=n$ and are, in particular, min-ready.
 Take $t\in R^{\prime}$ arbitrarily.
-As in the proof of Lemma 5.7, employing Lemma 5.11 instead of Lemma 5.3, one shows that $\bigcap_{i\leq n}\operatorname{\textnormal{minx}}(t_{i})\neq\varnothing.$ Let $t^{1},\dots,t^{n}$ denote all cyclic permutations of the tuple $t$.
-We have $t^{i}\in R$ for all $i\leq n$ by cyclicity of $R$.
-By injectivity of $\operatorname{\textnormal{lex}}^{[n]}$}, the tuple $s\coloneqq{\operatorname{\textnormal{lex}}^{[n]}}(t^{1},\dots,t^{n})$ is contained in $R^{\prime}$.
-Moreover, observe that $\min(s)={\operatorname{\textnormal{lex}}^{[m]}}(\min(t),\dots,\min(t))$ and $\operatorname{\textnormal{minx}}(s_{i})=\bigcap_{j\leq n}\operatorname{\textnormal{minx}}(t_{j})$ for all $i\leq n$.
+As in the proof of Lemma 6.5, employing Lemma 6.11 instead of Lemma 6.1, one shows that $\bigcap_{i\leq n}\operatorname{minx}(t_{i})\neq\varnothing.$ Let $t^{1},\dots,t^{n}$ denote all cyclic permutations of the tuple $t$.
+We have $t^{i}\in R^{\prime}$ for all $i\leq n$ by cyclicity of $R$ and Lemma 3.7.
+The tuple $s\coloneqq{\operatorname{lex}^{[n]}}(t^{1},\dots,t^{n})$ is contained in $R^{\prime}$.
+Moreover, observe that $\min(s)={\operatorname{lex}^{[m]}}(\min(t),\dots,\min(t))$ and $\operatorname{minx}(s_{i})=\bigcap_{j\leq n}\operatorname{minx}(t_{j})$ for all $i\leq n$.
 In particular, $s\in R^{\prime}$ is min-clean.